### v1

Abstract.
We survey theory developed over the past 10 years of semirings which need not be additively cancellative. The main feature is a specified “null ideal” $\mathcal{A}_{0}$ of a semiring $\mathcal{A},$ taking the place of a zero element, which permits generalizations of the classical algebraic theory to polynomials and their roots, algebraic geometry, matrices, linear algebra, varieties, categories, and module theory. The “pair” $(\mathcal{A},\mathcal{A}_{0})$ is studied along the lines of universal algebra.
Key words and phrases:
bimagma, dependence, function semiring, hyperfield, hyperpair, Krasner, linear algebra, magma, matrix, metatangible, pair, Property N, polynomial, residue semiring, subgroup, semiring, quotient semiring, supertropical, root, tensor, vector space.
2020 Mathematics Subject Classification:
Primary 16Y20, 16Y60, ; Secondary 12K10, 14T10, 06A12, 20M25.
Thanks to Tomer Bauer for helpful comments
1. Overview
A semigroup is a set with an associative operation, often denoted as concatenation. Occasionally, we use partial semigroups which mean that the operation need not be defined on every pair of elements, but $(a_{1}a_{2})a_{3}=a_{1}(a_{2}a_{3})$ whenever either side is defined.
An additive semigroup is a commutative semigroup, with the operation denoted by “++,” endowed with a neutral element $\mathbf{0}.$ The trivial additive semigroup is $\{\mathbf{0}\}$, and the two additive semigroups of order 2 are $\mathbb{Z}_{2}$ and the Boolean semigroup $(\mathbb{B},+)$, where $1+1=1$. An additive semigroup $\mathcal{A}$ is idempotent (also called a band) if $a+a=a$ for all $a\in\mathcal{A}.$ Thus, $(\mathbb{B},+)$ is idempotent.
By monoid we mean a semigroup under multiplication, containing $\mathbf{1}.$ Likewise for partial monoid. There is a considerable literature on semigroups, inspired in part by the example of the multiplicative monoid of a ring. If we want to permit two operations we are led to the following definition.
[…]
Semiring theory starts with the natural numbers $\mathbb{N}$, and semirings were introduced formally by Vandiver [55]. Golan’s book [16] is still an excellent resource, and [16, Chapter 1] has a wealth of examples. Semirings are a useful structure in computer programming, cf. [37] since one proceeds step by step.
There is a thriving theory of semigroups and monoids, but much less so with semirings. A satisfactory categorical theory for monoids is provided in [10], but it seems that the presence of the extra operation of addition without negation often complicates the situation and hampers basic categorical results.
Semirings can be studied as $\Omega$-algebras in universal algebra [21, Chapter 2]. In particular, one can define the direct product of semirings $\prod_{i\in I}\mathcal{A}_{i}.$ One calls two structures isomorphic (written $\cong$) when there is a 1:1 correspondence which respects their operations and distinguished elements (in the case of semirings, $\cdot,\mathbf{0},\mathbf{1}$).
However, when one wants to dig deeper, lack of negation makes it difficult to utilize roots of polynomials and determinants of matrices, two of the most important tools of algebras.
Another important area of ring theory is through the study of modules. Although one can define projective and injective modules over semirings, categorically without difficulty, it is much harder to obtain basic results in the homology theory of modules over semirings.
Since the theory of semirings has lagged far behind ring theory, researchers have compensated by considering semifields [56, 57, 20], which are semirings $\mathcal{A}$ for which $(\mathcal{A}\setminus\{\mathbf{0}\},\cdot)$ is a group; then group theory can be applied to the multiplicative structure instead of the additive structure. However this is rather restrictive when one aims for a general structure theory of semirings.
This article describes an attempt over the past 20 years to find a general algebraic framework developed for semirings and related structures, such as hyperfields, often used in conjunction with tropical algebra/geometry. The initial idea can be found in [12, 14], which was formulated as supertropical algebra in [23, 26] and was implemented for blueprints in [38, 39], and put in a more general context in 2016 in [50], in terms of a “negation map.” More recently this approach was broadened further in [33, 4] to a “pair” satisfying Property N, which still is sufficient often to provide a robust structure theory, and we shall study such pairs in this paper. Examples treated in §2.1.3 include supertropical pairs, tropicalized pairs, hyperpairs, function pairs, polynomial pairs, and matrix pairs. Categorically we shall see that actually there are three different types of morphisms, each giving rise to its own structure theory.
While studying semirings, we have encountered a fascinating connection between multiplication and addition, where the semiring has an underlying multiplicative monoid on which the additive structure acts, and this will drive forward most of the theory presented here.
Some of the paper is expository, and it will be noted where considerable work remains to be done. Other parts are largely new.
1.1. Additively cancellative semirings
The semiring $\mathbb{N}$ is one of the most fundamental mathematical entities. Of course $\mathbb{N}$ can be embedded into the ring $\mathbb{Z}$ via Grothendieck’s celebrated construction. Nevertheless, many famous assertions in the literature are formulated for $\mathbb{N}$ rather than $\mathbb{Z},$ for example Lagrange’s theorem that every natural number is the sum of four squares, and Waring’s problem in which $7$ is written as a sum of seven cubes ($1+\dots+1)$ rather than two cubes ($2^{3}+(-1)^{3}).$
The semiring $\mathbb{N}$ can be generalized in several directions, as the set of:
  1. (1)
    positive elements of an ordered ring;
  2. (2)
    sums of squares in a commutative ring;
  3. (3)
    sums of $m$-powers in a commutative ring, for any $m;$
  4. (4)
    $p$-powers in a commutative ring of characteristic $p$, for any prime $p.$
MAJOR NOTE 1.1.
In each of these examples, the semiring is additively cancellative ($a+b_{1}=a+b_{2}$ implies $b_{1}=b_{2}$); this property is inherited from the overlying ring. One thereby is led to try to study an arbitrary additively cancellative semiring by embedding it into a ring, first embedding the additive semigroup into a group and then naturally extending multiplication when possible. In this paper, our emphasis will be on non-additively cancellative semirings.
For technical reasons to become clear later, we weaken the notion of semiring:
Definition 1.2.
An nd-semiring is an additive semigroup which also is a monoid under multiplication, but multiplication need not distribute over addition.
1.2. Idempotent semirings
Example 1.3.
Vandiver’s example [55, p. 516], denoted $\mathbb{V}_{n}$, in which $\mathbb{N}$ is truncated at some number $n,$ identifying $m$ with $n$ for all $m\geq n.$ Additive cancellation clearly fails, since $n+1=n+0=n.$
A more sophisticated example is the set of finitely generated modules over a ring, where the sum is the direct sum, and the product is the tensor product.
But we are mainly interested in the other extreme. We say that a semiring $\mathcal{A}$ is idempotent if the semigroup $(\mathcal{A},+)$ is idempotent.
For example, the power set $(\mathbf{P}(S),\cup,\cap,\emptyset,S)$ of a set $S$, is an idempotent semiring. (Note that one could reverse $\cup$ and $\cap$ and still have an idempotent semiring!) Similarly, the set of ideals of a ring is an idempotent semiring.
Remark 1.4.
  1. (i)
    An idempotent semiring cannot be additively cancellative, since $a+\mathbf{0}=a=a+a.$
  2. (ii)
    Every idempotent semiring is zero sum free (ZSF) [16, p. 4], in the sense that $a_{1}+a_{2}=\mathbf{0}$ implies $a_{1}=a_{2}=\mathbf{0}$, since $a_{1}=a_{1}+(a_{1}+a_{2})=a_{1}+a_{2}=\mathbf{0}.$ [^2] ZSF semirings are called “antirings” in [11] and “lacking zero sums” in [25].
1.2.1. The max-plus algebra
In the max-plus algebra, the sum of two numbers is taken to be their maximum, and the product is taken to be their original sum. The “max-plus” algebra is bipotent in the sense that $a+b=b$ whenever $a\leq b$.
More generally, any ordered abelian monoid $(\mathcal{G},+)$ gives rise to a bipotent semiring $\mathcal{G}\cup\{\mathbf{0}\}$, with multiplication taken to be the original addition of $\mathcal{G},$ and addition taken to be the maximum ($\mathbf{0}$ formally taken to be the minimal element, which absorbs in multiplication).
The most basic example is the Boolean semifield $\mathbb{B}:=\mathbb{V}_{2}=\{0,1\}$ (taking the trivial monoid $\mathcal{G}=\{1\}$), contained canonically in every idempotent semiring.
1.2.2. Connection to tropical geometry
As explained in [15], the max-plus algebra is the foundation of tropical geometry, obtained via a limiting process of other semirings. Namely, for any real number $m,$ define a semiring structure on $\mathbb{R}$ by $a\bigodot_{m}b=a+b$ and $a\bigoplus_{m}b=m^{\log_{m}a+\log_{m}b}.$ [^3] A slightly different approach is to take $a\bigoplus_{m}b$ to be $\root m \of{a^{m}+b^{m}}$. Then $\lim_{m\to\infty}a\bigoplus_{m}b=\max\{a,b\}.$
The connection to tropical geometry is as follows: Given an algebraic curve over $\mathbb{C},$ one obtains an “amoeba” by taking the logarithm of the absolute values of the coordinates to the base $m,$ and the limiting case for $m\to\infty$ looks like a stick figure, called a “tropical curve.” Tropicalization greatly simplifies the geometry, although preserving key enumerative properties, and tropical mathematics has become a very active subject, as seen in [22, 40]. An excellent introduction to algebraic aspects of semirings in tropical theory is found in [5].
With the advent of the accompanying F1-algebra/geometry, cf. [9], researchers have been led to tackle the theory of idempotent semirings (and semifields).
MAJOR NOTE 1.5.
This revolutionary idea has several serious drawbacks:
  • In order to use logarithms, one has to pass through $\mathbb{R}$, which is not algebraically closed.
  • The limit procedure of having the base of logarithms go to infinity, although conceptually clear, is difficult for computations.
  • The lack of negatives hampers the algebraic development of the theory, for example in linear algebra.
1.3. Valuations and the supertropical semiring
The first two deficiencies of Note 1.5 are overcome by taking the field $\mathbb{K}$ of Newton-Puiseux series $f=\sum_{\alpha\in\mathbb{Q}}f_{\alpha}t^{\alpha}$, where the coefficients of $f_{\alpha}$ are in $\mathbb{C}$ (or any algebraically closed field) and the exponents of $t$ in of $f_{\alpha}$ are all in $\frac{1}{m}\mathbb{Z}$ for suitable $m$ depending on $f.$ The field $\mathbb{K}$ is known to be algebraically closed, so in mathematical logic is elementarily equivalent to $\mathbb{C}.$ At the same time we have the Puiseux series valuation $v:\mathbb{K}\to\mathbb{Q}$ taking $f$ to $\alpha$, where $\alpha$ is the lowest exponent of a nonzero summand $f_{\alpha}t^{\alpha}$ of $f.$ We recall that a (nonarchimedian) valuation from an integral domain $K$ to an ordered abelian group $(\mathcal{G},+\,)$ is a multiplicative monoid homomorphism $v:K\setminus\{0\}\to\mathcal{G}$, i.e., with $v(ab)=v(a)+v(b),$ and satisfying the property $v(a+b)\geq\min\{v(a),v(b)\}$ for all $a,b\in K.$ We formally put $v(0)=\infty.$ Information about a valuation $v:K\to\mathcal{G}\cup\{\infty\}$ can be garnered from the value group $v(K)$, which can replace $\mathcal{G}$.
A well-known fact about valuations: If $v(a_{1}),\dots,v(a_{m})$ are distinct, then $v(\sum_{i=1}^{m}a_{i})=\min\{v(a_{i}):1\leq i\leq m\}.$ Thus, v-v corresponds to $\max$ when $v(a_{1}),\dots,v(a_{m})$ are distinct. For example, if $f=3t^{2}+9t^{4}$ and $g=5t^{3}+7t^{4}$ then $f+g=3t^{2}+5t^{3}+16t^{4}$ and $v(f+g)=-2=\max\{-2,-3\}=\max\{-v(f),-v(g)\}.$ But if $f=3t^{2}+9t^{4}$ and $g=-3t^{2}+7t^{3}+7t^{4}$ then $f+g=7t^{3}+16t^{4}$ and v(f+g)=3<2.-v(f+g)=-3<-2. In general, we do not know a priori $v(f+g)$ when $v(f)=v(g).$ This led Zur Izhakian [23] to the following notion:
1.3.1. The supertropical semiring
Suppose $\mathcal{G}$ is an ordered abelian monoid with absorbing minimal element $\mathbf{0}_{\mathcal{G}}$, and $\mathcal{T}$ is an abelian monoid, together with a monoid homomorphism $\mu:\mathcal{T}\to\mathcal{G}$. Take the action $\mathcal{T}\times\mathcal{G}\to\mathcal{G}$ defined by $a\cdot g=\mu(a)g.$ Then $\mathcal{A}:=\mathcal{T}\cup\mathcal{G}$ becomes an abelian monoid when we extend the given multiplications on $\mathcal{T}$ and on $\mathcal{G},$ also using the above action of $\mathcal{T}$ on $\mathcal{G}$.
Setting $\mu(g)=g$ for all $g\in\mathcal{G},$ we define addition on $\mathcal{A}$ by
$b_{1}+b_{2}=\begin{cases}b_{1}\text{ if }\mu(b_{1})>\mu(b_{2}),\\ b_{2}\text{ if }\mu(b_{1})<\mu(b_{2}),\\ \mu(b_{1})\text{ if }\mu(b_{1})=\mu(b_{2}).\end{cases}.$
We call $\mathcal{A}$ the supertropical semiring, denoted $\mathbb{T}(\mathcal{T};\mathcal{G})$. $\mathcal{A}$ is a semiring, but is not idempotent since $\mathbf{1}+\mathbf{1}=\mu(\mathbf{1})\neq\mathbf{1}.$ We shall deal with this phenomenon shortly. Also, $\mathcal{A}$ not a semifield since $\mathcal{G}$ is a proper ideal. Note that $\mathcal{A}$ contains the supertropical Boolean semiring $\mathcal{B}=\{\mathbf{0},\mathbf{1},\mu(\mathbf{1})\},$ where $\mathbf{1}+\mathbf{1}=\mu(\mathbf{1}).$
1.4. Tropical extensions
Occasionally it is convenient to refine the supertropical semiring even further.
Definition 1.6.
[“Tropical extensions” [2, Proposition 2.12], a special case of $\mathcal{L}$-layered semirings” in [4, §7.2] and [24]] We are given a semigroup $(\mathcal{L},+,\mathbf{0})$ and an ordered set $\mathcal{G}$. The tropical extension $\mathcal{L}\rtimes\mathcal{G}$ of $\mathcal{L}$ consists of the set $\mathcal{L}\times\mathcal{G}$, made into a semigroup with the following addition:
[…]
When $\mathcal{L}$ is a semiring (resp. nd-semiring, cf. Definition 1.2) and $\mathcal{G}$ is an abelian ordered semigroup, we make $\mathcal{L}\rtimes\mathcal{G}$ into a semiring (resp. nd-semiring) under the componentwise operation $(\ell_{1},b_{1})(\ell_{2},b_{2})=(\ell_{1}\ell_{2},b_{1}+b_{2})$.
Remark 1.7.
$\mathcal{L}$ often is taken to be the Vandiver monoid $\mathbb{V}_{n}$ of Example 1.3. For $n=2,$ and $\mathcal{G}$ the max-plus algebra, there is an isomorphism from $\mathcal{L}\rtimes\mathcal{G}$ onto the supertropical semiring $\mathbb{T}(\mathcal{G};\mathcal{G})$.
1.5. Graded semirings
Definition 1.8.
A semiring $\mathcal{A}$ is graded by a semigroup $M$ if we can write $\mathcal{A}=\prod_{x\in M}A_{x}$ where $A_{x}A_{y}\subseteq A_{x+y}$. Some cases of especial note:
  • (The “supercase”) $\mathcal{A}$ is graded by the group $(\mathbb{Z}_{2},+)$.
  • (The “supertropical” case) $\mathcal{A}$ is graded by the Boolean monoid $(\mathbb{B},+)$.
  • (The “layered” case) $\mathcal{A}$ is graded by $\mathbb{V}_{n}$.
Each of these cases arises prominently, but an algebraic theory has been developed thoroughly only in the supertropical case.
1.6. Hyperrings and Hyperfields
Viro [58] (anticipated by [41, 44]) found another way of viewing addition for the supertropical semifield. Instead of $a+a=\mu(a)$, he interpreted $a+a$ as the multivalued sum $\{b:b\leq a\}.$ So let us formalize multivalued sums. Define $\mathcal{P}^{\sharp}(\mathcal{H}):=\mathcal{P}(\mathcal{H})\setminus\emptyset.$ We follow [36, 58], as extended by [3].
Definition 1.9.
Let $\mathcal{H}$ be a set.
  1. (1)
    We are given a commutative multivalued addition $\boxplus:\mathcal{H}\times\mathcal{H}\to\mathcal{P}^{\sharp}(\mathcal{H}),$ which is associative in the sense that if we define
    $a\boxplus S=S\boxplus a=\bigcup_{s\in S}\ a\boxplus s,$
    then $(a_{1}\boxplus a_{2})\boxplus a_{3}=a_{1}\boxplus(a_{2}\boxplus a_{3})$ for all $a_{i}$ in $\mathcal{H}.$
  2. (2)
    $\mathcal{H}$ is viewed as the set of singletons in $\mathcal{P}^{\sharp}(\mathcal{H}),$ identifying $a\in\mathcal{H}$ with $\{a\}$.
  3. (3)
    $(\mathcal{H},\boxplus,\mathcal{\mathbf{0}_{\mathbf{H}}})$ is a hypersemigroup when $\mathcal{H}$ has an element $\mathbf{0}_{\mathbf{H}}$, satisfying $\mathbf{0}_{\mathbf{H}}\boxplus a=a\boxplus\mathbf{0}_{\mathbf{H}}=a$ for all $a\in\mathcal{P}^{\sharp}(\mathcal{H})$.
  4. (4)
    A hypergroup is a hypersemigroup for which every element $a\in\mathcal{H}$ has a unique hypernegative $a\in\mathcal{H}$, in the sense that $\mathcal{\mathbf{0}_{\mathbf{H}}}\in a\boxplus(-a).$
Definition 1.10.
  1. (1)
    A hypersemiring (resp. hyperring) is a hypersemigroup (resp. hypergroup) $\mathcal{H}$, which also is a multiplicative monoid, $(\mathcal{H},\cdot)$, with $\mathbf{0}_{\mathbf{H}}$ absorbing and $\cdot$ distributing over hyperaddition.
  2. (2)
    A hyperfield is a hyperring $\mathcal{H}$ for which $(\mathcal{H}\setminus\{\mathbf{0}_{\mathbf{H}}\},\cdot)$ is a group.
Lemma 1.11.
  1. (i)
    Any hypersemigroup $(\mathcal{H},\boxplus,\mathbf{0}_{\mathbf{H}})$ induces a semigroup $(\mathcal{P}^{\sharp}(\mathcal{H}),\boxplus,\{\mathbf{0}_{\mathbf{H}}\}),$ with addition given by
    $S_{1}\boxplus S_{2}=\cup\{s_{1}\boxplus s_{2}:s_{i}\in S_{i}\}.$
  2. (ii)
    When $\mathcal{H}$ is a hypersemiring, then $(\mathcal{P}^{\sharp}(\mathcal{H}),\cdot,\{\mathbf{1}\})$ is a monoid, where we define $S_{1}\cdot S_{2}=\{a_{1}a_{2}:a_{i}\in S_{i}\}$, which satisfies “single distributivity”:
    (1.2) $a\cdot\boxplus S_{i}=\boxplus(a\cdot S_{i});\quad(\boxplus S_{i})\cdot a=\boxplus(S_{i}\cdot a),\qquad\forall a\in\mathcal{H},\ S_{i}\in\mathcal{P}^{\sharp}(\mathcal{H}).$
Proof.   Associativity and single distributivity are checked elementwise.
Warning: Even when $a_{1},a_{2}\in a_{1}\boxplus a_{2},$ the hyperset $S\cdot(a_{1}\boxplus a_{2})$ could be properly contained in $S\cdot a_{1}\,\boxplus\,S\cdot a_{2},$ as we see in Example 1.12(iii).
1.6.1. Examples of hyperfields
We recall some major examples of hyperfields from [3, 36, 58]. A large collection of hyperrings is given in [43, §2], and [18], and related examples are in [45].
Example 1.12.
  1. (i)
    Our motivation came from the tropical hyperfield [44, 58] which consists of the set $\mathcal{H}=\mathbb{R}\cup\{-\infty\}$, with -$\infty$ as the zero element $\mathbf{0}_{\mathbf{H}}$ and 0 as the unit element $\mathbf{1}$, equipped with addition $a\boxplus b=\{a\}$ if $a>b$, $a\boxplus b=\{b\}$ if $a<b$, and $a\boxplus a=[-\infty,a]$. It is easy to see that there is an embedding from the supertropical semifield to the tropical hyperfield, given by $a\mapsto a$ and $\mu(a)\mapsto[-\infty,a].$
    A special case is the Krasner hyperfield [36] $\mathcal{K}:=\{0,1\}$, with the usual multiplication law, and with hyperaddition defined by $x\boxplus 0=0\boxplus x=x$ for all $x,$ and $1\boxplus 1=\{0,1\}$. $\mathcal{P}^{\sharp}(\mathcal{K})=\mathcal{K}\cup\{\mathcal{K}\}$ is isomorphic to $\mathcal{B}$, by the map $0\mapsto\mathbf{0},\ 1\mapsto\mathbf{1},\ \{0,1\}\mapsto\mu(\mathbf{1}).$
  2. (ii)
    The hyperfield of signs $\mathcal{S}:=\{1,0,-1\}$, with the usual multiplication law, and hyperaddition defined by $\boxplus 1=1, 11$$\boxplus-1=-1$, $x\boxplus 0=0\boxplus x=x$ for all $x,$ and $1\boxplus-1=-1\boxplus 1=\{0,1,-1\}$. $\mathcal{S}\cup\{\mathcal{S}\}$ is a semiring isomorphic to $\widehat{\mathbb{B}}$, by sending $0\mapsto\mathbf{0},$ $1\mapsto(1,0)$, $1\mapsto(0,1)$, and $\{0,1,-1\}\mapsto(1,1).$
  3. (iii)
    The phase hyperfield. Let $S^{1}$ denote the complex unit circle, and take $\mathcal{H}=S^{1}\cup\{0\}$. Nonzero points $a$ and $b$ are antipodes if a=b.a=-b. Multiplication is defined as usual (corresponding on $S^{1}$ to addition of angles). We denote (ab)(a\,b) for an open arc of less than 180 degrees connecting two non-antipodal points $a,b$ of the unit circle. The hypersum is given, for $a,b\neq 0$, by
[…]
The phase hyperfield fails double distributivity, since, for $a\neq-b,$ $(a\,b)\boxplus(a\,b)=\mathcal{H}$ whereas $\{-1,0,1\}(a\,b)$ is $(a\,b)\cup\{0\}\cup-(a\,b)$.
1.6.2. Krasner’s residue construction
There also is a general construction of Krasner [36], done more generally in [51].
Theorem 1.13.
[4, Theorem D]
  1. (i)
    Suppose $(\mathcal{S},+)$ is a semigroup and $f:\mathcal{S}\to\bar{\mathcal{S}}$ is any set-theoretic map onto a set $\bar{\mathcal{S}}$. Define hyperaddition $\boxplus:\overline{\mathcal{S}}\to\mathcal{P}(\overline{\mathcal{S}})$ by $\overline{a}\boxplus\overline{a^{\prime}}=\{\overline{a+a^{\prime}}:f(a)=\overline{a},f(a^{\prime})=\overline{a^{\prime}}\},$ for $a,a^{\prime}\in{\mathcal{S}}.$ Then $\overline{\mathcal{S}}$ is a hypersemigroup (i.e., is associative). If $\mathcal{S}$} is a group then $\overline{\mathcal{S}}$ is a hypergroup, where $\overline{a}=\overline{-a}.$
  2. (ii)
    Suppose a monoid $G$ acts on a semiring $(R,+)$, such that the orbits are multiplicative, i.e., $(g_{1}a)(g_{2}a^{\prime})=g_{1}g_{2}(aa^{\prime}).$ Write $\overline{a}=Ga,$ for $a\in R.$ Then $\overline{R}:=R/G$ is a hypersemiring under the hyperaddition of (i), with $\overline{a}\overline{a^{\prime}}:=\overline{aa^{\prime}}.$
Proof.   (i) As in [36], given in [33, Proposition 2.13]. Namely
$(\overline{a}\boxplus\overline{a^{\prime}})+\boxplus\overline{a^{\prime\prime}}=\{\overline{a+a^{\prime}+a^{\prime\prime}}:f(a)=\overline{a},f(a^{\prime})=\overline{a^{\prime}},f(a^{\prime\prime})=\overline{a^{\prime\prime}}\}=\overline{a}\boxplus(\overline{a^{\prime}}\boxplus\overline{a^{\prime\prime}}).$
(ii) First note by induction that $\boxplus b_{i}\mathcal{G}=\{\sum a_{i}\mathcal{G}:\quad a_{i}\in b_{i}\mathcal{G}\}.$ Then
$a\mathcal{G}(\boxplus_{i=1}^{m}b_{i}\mathcal{G})=\left\{a\mathcal{G}\sum a_{i}^{\prime}\mathcal{G}:\ a_{i}^{\prime}\in b_{i}\mathcal{G}\right\}=\left\{(\sum aa_{i}^{\prime})\mathcal{G}:\,a_{i}^{\prime}\in b_{i}\mathcal{G}\right\}=\boxplus_{i=1}^{m}a\mathcal{G}b_{i}\mathcal{G},$
implying $\overline{R}$ is a hypersemiring.
Example 1.14.
The major example of (ii) is when $G$ is a multiplicative subgroup of $R$. If $R$ is a field, then $R/G$ is a hyperfield, known in the literature as the quotient hyperfield.
1.6.3. Description of major hyperfields as quotient hyperfields
We want to identify the different examples of Example 1.12 as quotient hyperfields. Towards this end, an isomorphism $f:\mathcal{H}\to\mathcal{H}^{\prime}$ of hyperfields is defined as a group isomorphism satisfying $f(a_{1}\boxplus a_{2})=f(a_{1})\boxplus f(a_{2})$ for all $a_{i}\in\mathcal{H}.$
Example 1.15.
Here are isomorphisms from Examples 1.12 to quotient hyperfields.
  1. (i)
    The tropical hyperfield is isomorphic to the quotient hyperfield $F/G$, where $F$ denotes any field with a surjective valuation $v:F\to\mathbb{R}\cup\{+\infty\}$, and $G:=\{f\in F:v(f)=0\}$.
  2. (ii)
    The Krasner hyperfield is isomorphic to the quotient hyperfield $F/F^{\times}$, for any field $F$.
  3. (iii)
    The hyperfield of signs is isomorphic to the quotient hyperfield $K/K_{>0}$ for each linearly ordered field $(K,\leq)$, where $K_{>0}$ is the group of positive elements of $K$.
  4. (iv)
    The phase hyperfield is isomorphic to the quotient hyperfield $\mathbb{C}/\mathbb{R}_{>0}$.
Remark 1.16.
Suppose that the hypernegative $\mathbf{1}$ exists in the quotient hypermonoid $\mathcal{H}=\mathcal{M}/G$. Define $e=\mathbf{1}\boxplus(-\mathbf{1})\in\mathcal{P}^{\sharp}(\mathcal{H})$ (so $\mathbf{0}_{\mathbf{H}}\in e$).
If $G=\{\pm 1\}$, then $\mathbf{0}_{\mathbf{H}}\in a\boxplus a$ for all $a\in\mathcal{H}.$
The failure of distributivity in some quotient hyperfields enhances the following result.
Lemma 1.17.
[51, Lemma 3.5] If $1\in G$, then $ee=e\boxplus e$ in $\mathcal{M}/G$.
It was thought until quite recently that non-quotient hyperfields are rare, but large classes of nonquotient hyperfields have been found in [42, 43, 18, 19].
1.7. Semiring constructions
Many standard constructions in algebra are applicable to semirings.
1.7.1. Doubling
One way of circumventing the lack of negatives is by doubling a semiring $\mathcal{A}$, inspired by Grothendieck’s familiar construction of $\mathbb{Z}$ from $\mathbb{N}$:
Definition 1.18.
Define the doubled semiring $\widehat{\mathcal{A}}=\mathcal{A}\times\mathcal{A}.$ Twist multiplication on $\widehat{\mathcal{A}}$ is defined as follows:
(1.3) $(b_{1},b_{2})\cdot_{\operatorname{tw}}(b^{\prime}_{1},b^{\prime}_{2})=(b_{1}b^{\prime}_{1}+b_{2}b^{\prime}_{2},b_{1}b^{\prime}_{2}+b_{2}b^{\prime}_{1}),\qquad(b_{1},b_{2}),(b^{\prime}_{1},b^{\prime}_{2})\in\widehat{\mathcal{A}}$.
Remark 1.19.
The doubled semiring $\widehat{\mathcal{A}}$ is $(\mathbb{Z}_{2},+)$-graded, the 0 component being $\mathcal{A}\times\{\mathbf{0}\},$ and the $1$ component being $\{\mathbf{0}\}\times\mathcal{A}.$ Also $\widehat{\mathcal{A}}$ has an equivalence relation given by $(b_{1},b_{2})\simeq(b^{\prime}_{1},b^{\prime}_{2})$ if $b_{1}b^{\prime}_{2}=b_{2}b^{\prime}_{1}$, but we do not apply it, since it would yield a ring. Instead Gaubert [14], who writes $\widehat{\mathbb{R}}$ as $\mathbb{S}$}, for “symmetrized,” viewed the first component $\mathcal{A}$ of $\widehat{\mathcal{A}}$, denoted $\mathcal{A}^{+},$ as the positive part, the second component, denoted $\mathcal{A}^{-},$ as the negative part, and the diagonal $\{(a,a):a\in A\}$ as the null part. $\mathcal{A}^{+}$ is isomorphic to $\mathcal{A}$ via $(a,0)\mapsto a$ and when $\mathcal{A}$} is idempotent, so is $\widehat{\mathcal{A}}$.
Remark 1.20.
The absolute value $(b_{1},b_{2})||:=b_{1}+b_{2}$ is a homomorphism $\widehat{\mathcal{A}}\to\mathcal{A}$, which is a retract of the injection of $\mathcal{A}$ to the positive part of $\widehat{\mathcal{A}}$.
1.7.2. Function semigroups and semirings, cf. [16, Chapter 2]
Definition 1.21.
Suppose $X$ is a set with a distinguished element $\mathbf{0}.$ Write $X^{S}$ for the functions from a set $S$ to $X$. For $f\in X^{S}$ define the support $\mathrm{supp}(f)=\{s\in S:f(s)\neq\mathbf{0}\},$ and $X^{(S)}$ for the functions with finite support.
Example 1.22.
Notation as above,
  1. (i)
    For an additive semigroup $\mathcal{A}$, $\mathcal{A}^{S}$ is an additive semigroup under elementwise addition of functions, whose zero element is the function with empty support, i.e., sending all elements to $\mathbf{0}$. The elements of $\mathcal{A}^{S}$ can be viewed as $S$-tuples, and the sub-semigroup $\mathcal{A}^{(S)}$ is the semigroup direct sum.
    (These are instances of the familiar direct product and direct sum of copies of $\mathcal{A}$ in universal algebra.)
  2. (ii)
    If $\mathcal{A}$ is a semiring and $S$ is considered an index set then $\mathcal{A}^{S}$ is a semiring with componentwise addition as in (i) and componentwise multiplication, where the unit element is the function sending all elements of $S$ to $\mathbf{1}.$
However, we are interested in a different semiring structure.
Lemma 1.23.
  1. (i)
    When $\mathcal{A}$ is a semiring and $S$ is an additive semigroup, $\mathcal{A}^{(S)}$ becomes a semiring under the addition of Example 1.22(i), but with convolution product $fg:s\mapsto\sum_{u+v=s}f(u)g(v).$ The unit element is the function sending $\mathbf{0}\mapsto\mathbf{1}$ and all other elements to $\mathbf{0}$.
  2. (ii)
    Generalizing (i), when $S$ is a partial monoid having the property that for any $s\in S$ there are only finitely many pairs $(u,v)\in S\times S:u+v=s,$ such as for $S$ well-ordered, then we can still define the convolution product.
1.7.3. Function polynomial semirings and related constructions
Example 1.24.
There is a host of instances of the function semiring construction over a semiring $\mathcal{A}$, under the convolution product.
  1. (i)
    We take a set of indeterminates $\Lambda=\{\lambda_{i}:i\in I\}$. The function polynomial semiring , is $\mathcal{A}^{({\mathbb{N}^{(I)}})}$where $f(\Lambda)$ denotes the function $(m_{i})\mapsto(f(m_{i})),$ for $(m_{i})\in\mathbb{N}^{(I)}.$ The monomials of $\mathcal{A}[\Lambda]$ are the polynomials with singleton support. Thus each polynomial is a sum of monomials. The case $\Lambda=\{\lambda\}$ is of special interest.
  2. (ii)
    The function power series semiring is $\mathcal{A}^{\mathbb{N}}$.
  3. (iii)
    The function Laurent series semiring is the set of functions $f\in\mathcal{A}^{\mathbb{Z}}$ for which $\mathrm{supp}(f)$ is bounded from below.
  4. (iv)
    The field $\mathbb{K}$ of Newton-Puiseux series from §1.3 is another instance.
1.7.4. Matrix semirings
The matrix semiring $M_{n}(\mathcal{A})$ can be defined as usual, with a set of matrix units $S=\{e_{i,j},1\leq i,j\leq n\}$ satisfying $e_{i,j}e_{k,\ell}=\delta_{j,k}e_{i,\ell}$ where $\delta$ denotes the Kronecker delta, and $\sum_{i=1}^{n}e_{i,i}=\mathbf{1}$. But matrix units are not closed under multiplication, so we need modify Lemma 1.23(iii) in order to view matrices as a function semiring. Namely, switching from addition on $S$ to multiplication to make the notation more familiar, suppose multiplication is defined only on a subset of $S$.
$\mathcal{A}^{S}$ becomes a semiring under the convolution product $fg:s\mapsto\sum_{uv=s}f(u)g(v).$ Take $\mathcal{A}^{S}$ where $S=\{(i,j):1\leq i,j\leq n\}$, under multiplication $(i,j)(k,\ell)=(i,\ell)$ when $j=k$, and is undefined otherwise; we get $M_{n}(\mathcal{A})$ by imposing the relation $\sum e_{i,i}=\mathbf{1}$.
1.7.5. Monoid semirings
Definition 1.25.
The monoid semiring $\mathcal{A}[\mathcal{M}]$ of a monoid $\mathcal{M}$ over a semiring $\mathcal{A}$ is the semigroup $\mathcal{A}^{(\mathcal{M})}$, under the addition of Example 1.22(i), which becomes a semiring with convolution product $fg=((\sum_{uv=s}f_{u}g_{v})s),$ where $f=(f_{u})$ and $g=(g_{v})$.
Example 1.26.
The polynomial semiring We write $\mathcal{A}[\Lambda]$ for the semiring $\mathcal{A}[\mathcal{M}]$, when $\mathcal{M}$ is the word monoid in $t$ commuting letters $\lambda_{1},\dots,\lambda_{t}$, and denoted $\mathcal{A}[\lambda]$ when $t=1.$ The semiring $\mathbb{N}[\lambda]$ plays a key role in the work of Lorscheid [38, 39] in $\mathbb{F}_{1}$-geometry.
MAJOR NOTE 1.27.
This product resembles Example 1.22(ii), and there is a canonical semiring homomorphism $\psi:\mathcal{A}[\mathcal{M}]\to\mathcal{A}^{(\mathcal{M})}$ sending $f=(f_{s})$ to the function $s\mapsto f_{s}$ in $\mathcal{A}.$ But $\psi$ need not be 1:1. For example, if $\mathcal{A}$ is the max-plus algebra then the polynomials $\lambda^{3}+\lambda^{2}+\lambda$ and $\lambda^{3}+\lambda$ define the same function! Actually, this situation also arises classically for polynomials over a field $F$ of $n$ elements; the polynomials $\lambda^{n}$ and $\lambda$ define the same function.
1.8. Modules and semirings over a monoid
Just as one studies rings as algebras over an underlying commutative ring (often a field), we should like to study semirings over an underlying commutative semiring, but it turns out that the key is its multiplicative structure, which is a monoid.
Definition 1.28.
$(\mathcal{T},\mathbf{1})$ is a monoid with a unit element $\mathbf{1}$. [^4] Strictly speaking, we should start with a partial monoid $tT$, to handle matrix units, and only require $\mathcal{T}\cup\{\mathbf{0}\}$ to be a monoid, but we leave those details to the reader.
  1. (1)
    A $\mathcal{T}$-module is an additive semigroup $(\mathcal{A},+,\mathbf{0}_{\mathcal{A}})$ together with a (left) $\mathcal{T}$-action $\mathcal{T}\times\mathcal{A}\to\mathcal{A}$ (denoted as concatenation), which is
    1. (a)
      associative, in the sense that $a_{1}(a_{2}b)=(a_{1}a_{2})b$ for all $a_{i}\in\mathcal{T},$ $b\in\mathcal{A}$, whenever $a_{1}a_{2}$ is defined, and $a_{1}(a_{2}b)$ is undefined whenever $a_{1}a_{2}$ is undefined.
    2. (b)
      zero absorbing, i.e. $a\mathbf{0}_{\mathcal{A}}=\mathbf{0}_{\mathcal{A}},\ \text{for all}\;a\in\mathcal{T}.$
    3. (c)
      distributive, in the sense that
      $a(b_{1}+b_{2})=ab_{1}+ab_{2},\quad\text{for all}\;a\in\mathcal{T},\;b_{i}\in\mathcal{A}.$
    4. (d)
      $\mathbf{1}b=b$ for all $b\in\mathcal{A}.$
  2. (2)
    We shall assume henceforth that $\mathcal{A}$ is a faithful $\mathcal{T}$-module, in the sense that if $ab_{1}=ab_{2}$, for $a\in\mathcal{T}$ and $b_{1},b_{2}\in\mathcal{A}$, then $b_{1}=b_{2}.$ Our convention is to write $a$ for an element of $\mathcal{T},$ and $b$ for an element of $\mathcal{A}.$ We call $\mathcal{T}$ the underlying monoid of tangible elements.
  3. (3)
    Define formally the monoid $\mathcal{T}_{0}=\mathcal{T}\cup\{\mathbf{0}\}$ with absorbing element $\mathbf{0}$, declaring $a_{1}a_{2}=\mathbf{0}$ in $\mathcal{T}_{0}$ whenever $a_{1}a_{2}$ is undefined in $\mathcal{T}$, and view $\mathcal{A}$ as a $\mathcal{T}_{0}$-module by declaring $\mathbf{0}a=\mathbf{0}_{\mathcal{A}}$ for all $a\in\mathcal{A}$.
  4. (4)
    A $\mathcal{T}$-module $\mathcal{A}$ is free if there is a base $\mathcal{B}$ such that every element of $\mathcal{A}$ can be written uniquely in the form $\sum a_{b}b:\ b\in\mathcal{B},\ a_{b}\in\mathcal{T}_{0}$.
Definition 1.29.
A $\mathcal{T}$-semiring is a semiring which is also a $\mathcal{T}_{0}$-module, such that the natural map $\iota:\mathcal{T}_{0}\to\mathcal{A}$ given by $\iota(a)=a\mathbf{1}_{\mathcal{A}}$ is an injection such that $\iota(\mathbf{1})=\mathbf{1}_{\mathcal{A}}$. (Thus, the $\mathcal{T}$-module action and multiplication by $\iota(\mathcal{T})$ within $\mathcal{A}$ are the same.) We use $\iota$ to view $\mathcal{T}_{0}\subseteq\mathcal{A}$. $\mathcal{A}$ is called admissible if $\mathcal{T}_{0}$ is a spanning subset of $(\mathcal{A},+).$
We also assume throughout this exposition that $ab=b\iota(a)$ and $b_{1}(ab_{2})=a(b_{1}b_{2})=(ab_{1})b_{2}$ for all $a\in\mathcal{T},$ $b_{i}\in\mathcal{A},$ i.e., the copy of $\mathcal{T}$ in $\mathcal{A}$ is central.
Example 1.30.
  1. (i)
    Suppose $\mathcal{A}$ is a $\mathcal{T}$-module.
    • For any set $S$, $\mathcal{A}^{S}$ is a $\mathcal{T}$-module under the pointwise action $(af)(s)=af(s).$
    • For any monoid $\mathcal{M}$, $\mathcal{A}[\mathcal{M}]$ is a $\mathcal{T}$-module under the action $a(f_{u})=(af_{u}).$
  2. (ii)
    The free $\mathcal{T}$-module is $\mathbb{N}^{(\mathcal{T})},$ viewing $\mathbb{N}$ as an additive semigroup.
MAJOR NOTE 1.31.
Many important properties are not preserved under the polynomial and matrix constructions. This is evident in hyperrings, for example. If $\mathcal{H}$ is a hyperring then $\mathcal{H}[\lambda]$ is not a hyperring in any natural way, since how would we define $a_{1}\lambda\boxplus a_{2}\lambda^{2}$ and $(\mathbf{1}\boxplus\lambda)^{2}$? Likewise, matrix constructions do not work well with hyperrings.
The difficulty is that we are not working directly with the hyperring, but rather its power set, which may fail to be distributive (cf. §2.2 below). Thus representation theory over hyperfields is problematic.
1.9. Homomorphisms
We consider $\mathcal{T}_{i}$-modules $\mathcal{M}_{i}$, and
Definition 1.32.
Suppose $\mathcal{A}_{i}$ are $\mathcal{T}_{i}$-modules, for $i=1,2.$
  1. (i)
    Our maps $f$ from $\mathcal{A}_{1}$ to $\mathcal{A}_{2}$ always will satisfy
    1. (a)
      $f(\mathcal{T}_{1})\subseteq\mathcal{T}_{2},$
    2. (b)
      $f(\mathbf{0}_{\mathcal{A}_{1}})=(\mathbf{0}_{\mathcal{A}_{2}}),$ and $f(\mathbf{1}_{\mathcal{A}_{1}})=\mathbf{1}_{\mathcal{A}_{2}},$
    3. (c)
      $f(ab)=f(a)f(b)$ for all $a\in\mathcal{T}_{1},\ b\in\mathcal{A}_{1}.$
  2. (ii)
    A $\mathcal{T}_{1}$-homomorphism $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a map satisfying
    $f(b_{1}+b_{2})=f(b_{1})+f(b_{2}),\quad\forall b_{i}\in\mathcal{A}_{1}.$
  3. (iii)
    $\operatorname{Hom}(\mathcal{A}_{1},\mathcal{A}_{2})$ is the set of $\mathcal{T}_{1}$-homomorphisms $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$.
Lemma 1.33.
Suppose $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a $\mathcal{T}_{1}$-homomorphism.
  1. (i)
    $\mathcal{A}_{2}$ is an $f(\mathcal{T}_{1})$-module, so we can replace $\mathcal{T}_{2}$ by $f(\mathcal{T}_{1}),$ and assume that the restriction $f_{\mathcal{T}_{1}}$ (defined $f_{\mathcal{T}_{1}}(a)=f(\iota(a))$ for $a\in\mathcal{T}_{1}$) is onto.
  2. (ii)
    Assuming $f_{\mathcal{T}_{1}}$ is onto, view $\mathcal{A}_{2}$ as a $\mathcal{T}_{1}$-module, by defining $ab_{2}=f(a)b_{2}$ for $a\in\mathcal{T}_{1},$ $b_{2}\in\mathcal{A}_{2}.$
  3. (iii)
    Let $\operatorname{id}:\mathcal{A}_{2}\to\mathcal{A}_{2}$ denote the identity, from $\mathcal{A}_{2}$ as $\mathcal{T}_{1}$-module to $\mathcal{A}_{2}$ as $\mathcal{T}_{2}$-module. Then $f$ is a composition of $\mathcal{T}_{1}$-homomorphisms $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ as $\mathcal{T}_{1}$-modules, satisfying $f(ab)=af(b)$ for $a\in\mathcal{T},$ followed by $\operatorname{id}.$
Proof.   (i) $f_{\mathcal{T}}(aa^{\prime})=f(\iota(aa^{\prime}))=f(\iota(a)f(\iota(a^{\prime})=f_{\mathcal{T}}(a)f_{\mathcal{T}}(a^{\prime}).$
(ii), (iii) A standard change of the underlying monoid. Note by definition that $af(b)=f(a)f(b)=f(ab).$
Remark 1.34.
Since Lemma 1.33 is canonical, we shall assume from now on, for all $\mathcal{T}_{1}$-homomorphisms $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$, that $\mathcal{T}_{1}=\mathcal{T}_{2}=\mathcal{T},$ and $f_{\mathcal{T}}$ always is the identity map. Thus a $\mathcal{T}$-homomorphism satisfies $f(ab)=af(b)$ for $a\in\mathcal{T},b\in\mathcal{A}_{1}$.
Lemma 1.35.
For $\mathcal{T}$-modules $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, $\operatorname{Hom}(\mathcal{A}_{1},\mathcal{A}_{2})$ is a $\mathcal{T}$-module under the operation $(f_{1}+f_{2})(b)=f_{1}(b)+f_{2}(b)$ and the action $af:b\mapsto af(b)$. $\operatorname{Hom}(\mathcal{A},\mathcal{A})$ is a $\mathcal{T}$-semiring under composition of homomorphisms, which we denote as $\mathrm{End}\mathcal{A}$.
Proof.   Standard.
Definition 1.36.
Suppose $A_{i}$ now are $\mathcal{T}$-semirings. A map $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is multiplicative if $f(b_{1}b_{2})=f(b_{1})f(b_{2}),$ $\forall b_{i}\in\mathcal{A}_{1}.$ A $\mathcal{T}$-semiring homomorphism $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a multiplicative map satisfying $f(b_{1}+b_{2})=f(b_{1})+f(b_{2}),$ $\forall b_{i}\in\mathcal{A}_{1}.$
MAJOR NOTE 1.37.
  • As explained in [21, Chapter 2], in universal algebra, homomorphisms are defined via congruences, where we recall that a congruence on $\mathcal{A}$ is an equivalence relation which viewed as a subset of $\mathcal{A}\times\mathcal{A}$ is a subalgebra. In particular, the trival congruence $\operatorname{diag}$ is $\{(a,a):a\in\mathcal{A}\}$. This approach fits in well with semigroups and semirings viewed via universal algebra, but does not always yield the categories that we desire, as we shall see, and other morphisms are described in §8.
  • Given a congruence $\operatorname{\Phi}$} on a semiring $\mathcal{A},$ one defines $\mathcal{A}/{\operatorname{\Phi}}$ as the semiring $\{[b]:b\in\mathcal{A}\}$, identifying $b_{1}]=[b_{2}]$ when $(b_{1},b_{2})\in{\operatorname{\Phi}}.$ (For example, $\mathcal{A}/\operatorname{diag}\cong\mathcal{A}.$)
  • If $f:\mathcal{A}_{1}\to\mathcal{A}_{2}$ is a homomorphism then for $b\in\mathcal{A}_{2},$ the inverse image $f^{-1}(b)$ is an ideal, not a congruence. This creates a certain tension, where sometimes we choose ideals and sometimes congruences, to be explained in Note 2.24.
2. Pairs
One can continue to make definitions parallel to rings, but one encounters obstacles when attempting to factor polynomials or to describe dependence of vectors, due to the lack of negation. This leads us to the objective of this paper, to find some extra structure which will support a plausible structure theory. The underlying idea, originating in Gaubert’s dissertation [14], is to replace $\mathbf{0}$, which is useless for LZS semirings, by a designated additive subsemigroup.
The groundwork has been laid for our main notion.
Definition 2.1.
We follow [4, 33].
  1. (1)
    A $\mathcal{T}$-pair $(\mathcal{A},\mathcal{A}_{0})$ is a $\mathcal{T}$-module $\mathcal{A}$, given together with a specified $\mathcal{T}$-submodule $\mathcal{A}_{0}$, i.e., $ab_{0}\in\mathcal{A}_{0}$ for all $a\in\mathcal{T}$ and $b_{0}\in\mathcal{A}_{0},$ and which also satisfies the converse:
    If $ab\in\mathcal{A}_{0}$ for $a\in\mathcal{T},$ then $b\in\mathcal{A}_{0}$ (which is obvious if $\mathcal{T}$ is a group).
  2. (2)
    From now on, excepting the Examples (iv),(v) of §2.1.3 and Definition mapr, the underlying monoid $\mathcal{T}$ is given ( until §8.1), and we simply call a $\mathcal{T}$-pair a “pair.” We call $\mathcal{A}_{0}$ the “null part” of $\mathcal{A}.$
MAJOR NOTE 2.2.
  1. (i)
    The notion of “pair” is general on purpose, in order to let $\mathcal{A}_{0}$ replace $\mathbf{0}$ in diverse algebraic situations. Following the spirit of universal algebra in [21], we define an abstract pair $(\mathcal{A},\mathcal{A}_{0})$ to consist of an $\Omega$-algebra with an action of a monoid $\mathcal{T}$ on $\mathcal{A}$ which preserves the universal algebra operations, and $\mathcal{A}_{0}$ is an $\Omega$-subalgebra closed under the $\mathcal{T}$-action.
    • For semiring pairs, we have Definition 2.14 below.
    • From this point of view, we do not need $\mathbf{0}\in\mathcal{A},$ and we could define $\mathcal{A}=\mathcal{T}\cup\{\infty\}$, where $a\infty=\infty$ for all $a\in\mathcal{T}.$ For example, take $\mathcal{T}=\{\mathbf{1}\}$ and $(\mathcal{A},\mathcal{A}_{0})=(\{\mathbf{1},\infty\},\{\infty\}).$
    • To define monoid pairs, we need the conditions $a(b_{1}b_{2})=(ab_{1})b_{2}$ and $a\mathcal{A}_{0}\subseteq\mathcal{A}_{0}$ for all $a\in\mathcal{T},$ $b_{i}\in\mathcal{A}.$
    • Lie pairs were treated in [13].
  2. (ii)
    An isomorphism of abstract pairs is an $\Omega$-algebra isomorphism which restricts to an $\Omega$-isomorphism of the null parts.
2.0.1. Surpassing relations and reversibility
We need some relation to replace equality. Lorscheid [38] used a symmetric relation; the one we adapt from [50], on the contrary, is antisymmetric on $\mathcal{T}.$
Definition 2.3.
  1. (i)
    A pre-order on a $\mathcal{T}$-module $\mathcal{A}$, denoted $\preceq$, is a set-theoretic pre-order that respects the $\mathcal{T}$-module structure, i.e., for all $b,b_{i}\in\mathcal{A}$:
    1. (a)
      For all $a\in\mathcal{T}$, $b_{1}\preceq b_{2}$ if and only if $ab_{1}\preceq ab_{2}.$
    2. (b)
      $b_{i}\preceq b_{i}^{\prime}$ implies $b_{1}+b_{2}\preceq b_{1}^{\prime}+b_{2}^{\prime}.$
  2. (ii)
    A pre-surpassing relation on a pair $(\mathcal{A},\mathcal{A}_{0})$, denoted $\preceq$, is a pre-order satisfying the condition
    • If $b\in\mathcal{A}_{0}$ then $\mathbf{0}\preceq b.$
  3. (iii)
    A surpassing relation on a pair $(\mathcal{A},\mathcal{A}_{0})$ is a pre-surpassing relation satisfying the condition
    • $a_{1}\preceq a_{2}$ for $a_{1},a_{2}\in\mathcal{T}_{0}$ implies $a_{1}=a_{2}.$
  4. (iv)
    We write $b_{1}\succeq b_{2}$ if $b_{2}\preceq b_{1}.$
  5. (v)
    For any pair $(\mathcal{A},\mathcal{A}_{0})$, a pre-surpassing relation $\preceq$ is
    1. (a)
      $\mathcal{T}$-reversible if $a_{0}^{\dagger}\preceq\sum_{i=1}^{n}a_{i}\in\mathcal{A}_{0}$ implies $a_{1}^{\dagger}\preceq a_{0}+\sum_{i=2}^{n}a_{i}\in\mathcal{A}_{0}$,
    2. (b)
      strongly $\mathcal{T}$-reversible if $a+b\in\mathcal{A}_{0}$ for $a\in\mathcal{T}$, implies $a^{\dagger}\preceq b$,
    for $a,a_{i}\in\mathcal{T}$, $b\in\mathcal{A}$.
Any pre-surpassing relation of a uniquely negated pair is a surpassing relation, by [4, Lemma 3.28].
Lemma 2.4.
If $b+b_{1}\preceq b_{2}$, then $b_{1}\preceq b^{\dagger}+b_{2}.$
Proof.   $b_{1}\preceq b_{1}+b^{\circ}=b^{\dagger}+(b+b_{1})\preceq b^{\dagger}+b_{2}$.
Lemma 2.5.
Given any strongly $\mathcal{T}$-reversible pre-surpassing relation $\preceq$ on  $(\mathcal{A},\mathcal{A}_{0})$,
  1. (i)
    $(\mathcal{A},\mathcal{A}_{0})$ is uniquely negated.
  2. (ii)
    $\preceq$ is a $\mathcal{T}$-reversible surpassing relation.
Proof.   (i) Suppose $a^{\prime}+a\in\mathcal{A}_{0}$. Then $\mathbf{0}^{\dagger}\preceq a^{\prime}+a,$ implying $a^{\prime}\preceq\mathbf{0}+a^{\dagger},$ so $a^{\dagger}=a^{\prime}$. By symmetry, $(a^{\prime})^{\dagger}=a,$ so $(a^{\dagger})^{\dagger}=a,$ and $\dagger$} is a negation map ()(-), and $a^{\prime}=(-)a$, so $(\mathcal{A},\mathcal{A}_{0})$ is uniquely negated.
(ii) If $a_{1}\preceq a_{2}$ then $a_{2}(-)a_{1}\succeq a_{1}(-)a_{1},$ implying $a_{2}=a_{1}$ by (i). Hence $\preceq$ is a surpassing relation.
Next, suppose $a_{0}^{\dagger}\preceq\sum_{i=1}^{n}a_{i}\in\mathcal{A}_{0}$ for $a_{i}\in\mathcal{T}$. Clearly $a_{0}+\sum_{i=1}^{n}a_{i}\in\mathcal{A}_{0}$, implying $a_{1}^{\dagger}\preceq a_{0}+\sum_{i=2}^{n}a_{i}\in\mathcal{A}_{0},$ by hypothesis.
Example 2.6.
  1. (i)
    Every pair $(\mathcal{A},\mathcal{A}_{0})$ has the “minimal” pre-surpassing relation $\preceq_{0}$, given by defining $b_{1}\preceq_{0}b_{2}$ when $b_{1}+b^{\prime}=b_{2}$ for some $b^{\prime}\in\mathcal{A}_{0}.$
  2. (ii)
    Power sets of hypersemirings all have the surpassing relation $\subseteq,$ which at times is more appropriate for hyperring theory than $\preceq_{0}$.
  3. (iii)
    $\succeq$ is also a pre-surpassing relation, and in particular so is $\succeq_{0}.$
  4. (iv)
    For a pair $(\mathcal{A},\mathcal{A}_{0})$ of the first kind, define the pre-surpassing relation $b_{1}\preceq^{\circ}b_{2}$ if $b_{1}^{\circ}+b_{2}^{\circ}=b_{2}^{\circ}.$
Gunn [17] recalls a property called fissure, that if $a_{0}\preceq\sum_{i=1}^{n}a_{i}\in\mathcal{A}_{0}$ for $a_{i}\in\mathcal{T}_{0}$, then there is $a\in\mathcal{T}_{0}$ such that $a_{0}\preceq a_{1}+a$ and $a\preceq\sum_{i=2}^{n}a_{i}.$
Lemma 2.7.
Suppose $(\mathcal{A},\mathcal{A}_{0})$ is uniquely negated and satisfies fissure. Then the surpassing relation $\preceq$ is strongly $\mathcal{T}$-reversible.
Proof.   Suppose $\mathbf{0}\preceq\sum_{i=1}^{t}a_{i}$ for $a_{i}\in\mathcal{T}$. Fissure implies there exists $a$ such that $\mathbf{0}\preceq a_{1}+a$ and $a\preceq\sum_{i=2}^{t}a_{i}$. But then $a=(-)a_{1}$, implying $(-)a_{1}\preceq\sum_{i=2}^{t}a_{i}$.
Definition 2.8.
Given a monoid pair $(\mathcal{A},\mathcal{A}_{0})$ (in the sense of Remark 2.2(i)) with a pre-surpassing relation $\preceq$, the $\preceq$-unit monoid $\mathcal{A}^{\preceq\times}$ is $\{b\in\mathcal{A}:\mathbf{1}\preceq b\}$, and $\mathcal{A}^{\preceq\times}_{0}=\mathcal{A}^{\preceq\times}\cap\mathcal{A}_{0}.$
Lemma 2.9.
$\mathcal{A}^{\preceq\times}$ is a monoid, and $(\mathcal{A}^{\preceq\times},\mathcal{A}_{0}^{\preceq\times})$ is a monoid pair.
Proof.   If $\mathbf{1}+c_{i}^{\circ}=b_{i}$ for $i=1,2$, then $\mathbf{1}+(c_{1}b_{2}+c_{2}b_{1}+c_{1}c_{2})^{\circ}=b_{1}b_{2}.$
2.0.2. Metatangible and weakly metatangible pairs
The following kinds of pairs are especially malleable.
Definition 2.10.
  1. (i)
    $(\mathcal{A},\mathcal{A}_{0})$ is weakly metatangible if $a_{1}+a_{2}\in\mathcal{T}\cup\mathcal{A}_{0}$ for any $a_{1},a_{2}$ in $\mathcal{T}.$
  2. (ii)
    ([4, Lemma 3.31] ) $(\mathcal{A},\mathcal{A}_{0})$ is metatangible if it is weakly metatangible and $\mathcal{A}=\mathcal{T}+\mathcal{A}_{0}$.
  3. (iii)
    As a special case, $(\mathcal{A},\mathcal{A}_{0})$ is $\mathcal{A}_{0}$-bipotent if $a_{1}+a_{2}\in\{a_{1},a_{2}\}\cup\mathcal{A}_{0}$ for any $a_{1},a_{2}$ in $\mathcal{T}.$
Theorem 2.11 ([4, Theorem A]).
For any metatangible pair $(\mathcal{A},\mathcal{A}_{0})$, every element $c$ of $\mathcal{A}$ has a “uniform presentation”, by which we mean there exists $c_{\mathcal{T}}\in\mathcal{T}$ and $m_{c}\in\mathbb{N}$, we have $c_{\mathcal{T}}\in\mathcal{T}$, such that either:
  1. (i)
    The pair $(\mathcal{A},\mathcal{A}_{0})$ is of the first kind, with $c=m_{c}c_{\mathcal{T}},$ or
  2. (ii)
    The pair $(\mathcal{A},\mathcal{A}_{0})$ is of the second kind, with one of the following:
    1. (a)
      $c=\mathbf{0}$, and $m_{c}=0.$
    2. (b)
      $c=c_{\mathcal{T}}\in\mathcal{T}$ and $m_{c}=1$, or
    3. (c)
      $c=c_{\mathcal{T}}^{\circ}$ and $m_{c}=2$.
In (i), for some $m_{0}\geq 0,$ $\mathcal{A}_{0}=\{m_{c}c_{\mathcal{T}}:m_{c}\geq m_{0}\text{ or $m$ is even}\}.$
In (ii), $\mathcal{A}=\{\mathbf{0}\}\cup\mathcal{T}\cup\mathcal{T}^{\circ}$, and $\mathcal{A}_{0}=\mathcal{T}_{0}^{\circ}$.
Lemma 2.12.
Suppose that a uniquely negated metatangible pair $(\mathcal{A},\mathcal{A}_{0})$ has a surpassing relation $\preceq$.
  1. (i)
    If $\sum a_{i}\in\mathcal{A}_{0}$ for $a_{i}\in\mathcal{T}$, then either $a_{1}=a_{2}=\dots=a_{t}$ or $(-)a_{1}\preceq\sum_{i=2}^{t}a_{i}.$
  2. (ii)
    $\preceq$ is $\mathcal{T}$-reversible.
  3. (iii)
    If $a_{1}+a_{2}+a_{3}\in\mathcal{A}_{0}$ for $a_{i}\in\mathcal{T}$, then either $a_{2}\preceq a_{3}(-)a_{1}$, or $a_{1}=a_{2}=(-)a_{3}$ with $a_{1}+a_{1}=a_{1}$, or $a_{1}=a_{2}=a_{3}$.
Proof.  (i) By induction on $t.$ If $t=2$ then $a_{1}+a_{2}\in\mathcal{A}_{0},$ so $a_{2}=(-)a_{1}.$ So may assume that $t\geq 3.$ If $a_{t-1}+a_{t}\in\mathcal{A}_{0},$ we can replace $a_{t-1}$ by $a_{t-1}+a_{t}$ and lower $t.$ Hence we may assume that $a_{t-1}+a_{t}\in\mathcal{A}_{0},$ and thus $a_{t-1}=(-)a_{t}.$ Permuting the indices shows $a_{t-2}=(-)a_{t}$ and $a_{t-2}=(-)a_{t-1}.$ Hence $a_{t-2}=(-)a_{t-2}$.
(ii) Suppose $a_{0}^{\prime}\preceq\sum_{i=1}^{t}a_{i},$ with $t\geq 3.$ Then $\sum_{i=0}^{t}a_{i}\succeq a_{0}+a_{0}^{\prime}$ so is in $\mathcal{A}_{0},$ and the assertion follow from (i).
(iii) We are given $\mathbf{0}\preceq a_{1}+a_{2}+a_{3}=a_{1}+(a_{2}+a_{3})$. By (i), we are done unless $(-)a_{2}\preceq a_{1}+a_{3}.$ If $a_{1}+a_{3}\in\mathcal{T}$ then $a_{1}+a_{3}=a_{2},$ so $a_{1}+a_{2}+a_{3}=a_{2}+a_{2}\in\mathcal{A}_{0},$ implying $a_{2}=(-)a_{2},$ so $a_{3}=(-)a_{3}$ and we are done.
Hence we assume that $a_{1}+a_{3}\in\mathcal{A}_{0}.$ Thus $a_{1}=(-)a_{3}.$ If $a_{1}+a_{2}\in\mathcal{T}$ then $a_{1}+a_{2}=(-)a_{3}=a_{1}.$ But $(-)a_{3}\preceq a_{1}+a_{2}=a_{2},$ implying $a_{2}=(-)a_{3}=a_{1}$.
MAJOR NOTE 2.13.
  1. (i)
[…]
  2. (ii)
    Metatangibility was assumed in much of [51] as a key ingredient of the theory. But in this paper, motivated by Lemma 2.12, we use the weaker hypothesis of $\mathcal{T}$-reversibility.
2.1. Semiring Pairs
Definition 2.14.
  1. (1)
    A semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is a pair for which $\mathcal{A}$ is a semiring.
  2. (2)
    A semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is said to be
    • proper if $\mathcal{A}_{0}\cap\mathcal{T}_{0}=\{\mathbf{0}\}$.
    • shallow if $\mathcal{T}\cup\mathcal{A}_{0}=\mathcal{A}.$
  3. (3)
    A semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is of the first kind if $\mathbf{1}+\mathbf{1}\in\mathcal{A}_{0}$ and of the second kind if $\mathbf{1}+\mathbf{1}\notin\mathcal{A}_{0}$ (And thus $a+a\notin\mathcal{A}_{0}$ for all $a\in\mathcal{T}$).
  4. (4)
    A semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is an $\mathcal{A}_{0}$-domain if $\mathcal{A}_{0}$ is a strongly prime ideal, i.e., $b_{1}b_{2}\in\mathcal{A}_{0}$ implies $b_{1}\in\mathcal{A}_{0}$ or $b_{2}\in\mathcal{A}_{0}$.
MAJOR NOTE 2.15.
“Classical algebra” is when $R$ is an algebra over a commutative ring $C$, and $\mathcal{T}_{0}=C$, viewed as a multiplicative monoid. Or we could simply take $\mathcal{T}=\{\mathbf{1}\}.$ But in this paper we are more interested in other situations.
The idea here is that in proper semiring pairs we “see” the elements $\mathcal{T},$ but $\mathcal{A}_{0}$ controls much of the structure. Historically, “first kind” is important since it gives Izhakian’s supertropical algebra “supertropical algebra” and the “flavor” of characteristic 2. “Second kind” is more closely related to classical algebra of characteristic $\neq 2$.
Lemma 2.16.
Every proper shallow pair is an $\mathcal{A}_{0}$-domain.
Proof.   If $b_{1},b_{2}\in\mathcal{A}\setminus\mathcal{A}_{0}$ then $b_{1},b_{2}\in\mathcal{T},$ so $b_{1}b_{2}\notin\mathcal{A}_{0}.$
2.1.1. Property N
A key issue is how $\mathcal{T}$ interfaces with $\mathcal{A}.$
Definition 2.17.
[4, §3.1] A pair $(\mathcal{A},\mathcal{A}_{0})$ satisfies Property N (for “negation”) if there is an invertible element $\mathbf{1}^{\dagger}\in\mathcal{T}$ such that, defining $b^{\dagger}=\mathbf{1}^{\dagger}b$ and $b^{\circ}=b+b^{\dagger}$:
  1. (i)
    $e:=\mathbf{1}+\mathbf{1}^{\dagger}\in\mathcal{A}_{0}$.
  2. (ii)
    (Uniqueness of $e$) If $\mathbf{1}+a\in\mathcal{A}_{0}$ for $a\in\mathcal{T},$ then $\mathbf{1}+a=e.$
  3. (iii)
    $b^{\circ}\in\mathcal{A}_{0}$ for each $b\in\mathcal{A}$.
(iii) is automatic for semiring pairs, and for $b\in\mathcal{T}$. When the pair $(\mathcal{A},\mathcal{A}_{0})$ is of the first kind, we may (and shall) take $\mathbf{1}^{\dagger}=\mathbf{1}.$
Lemma 2.18 ([4, Lemma 3.10]).
$\mathbf{1}^{\dagger}e=e.$ Thus, in a semiring pair, $e^{2}=e+e.$
MAJOR NOTE 2.19.
Any $\mathcal{T}$-module $\mathcal{A}$ gives rise to the trivial pair $(\mathcal{A},\mathbf{0}),$ but when $\mathcal{A}$ is idempotent, it lacks Property N, being ZSF, and this difficulty provided the motivation for the study of pairs.
  1. (i)
    From now on, every given pair $(\mathcal{A},\mathcal{A}_{0})$ also is assumed to satisfy Property N. We arbitrarily choose $\mathbf{1}^{\dagger}$, which need not be uniquely determined.
  2. (ii)
    In earlier incarnations, $\mathbf{1}^{\dagger}$ was required to be unique, designated $(-)\mathbf{1}$, with $((-)\mathbf{1})^{2}=\mathbf{1},$ yielding a unique “negation map” $b\mapsto((-)\mathbf{1})b.$ In this case we write ()b(-)b for $b^{\dagger},$ and say that the pair $(\mathcal{A},\mathcal{A}_{0})$ is uniquely negated. Then we write $b_{1}(-)b_{2}$ for $b_{1}+((-)b_{2}).$
    However, one of the themes of this exposition is that the uniqueness of $e$ often is all that is needed.
Examples are given in [4, §5], and below in §2.1.3.
2.1.2. Paired maps and homomorphisms
Definition 2.20.
Suppose $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ are pairs.
  1. (i)
    A paired map $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a map $f:\mathcal{A}\to\mathcal{A}^{\prime}$ satisfying $f(\mathcal{A}_{0})\subseteq\mathcal{A}_{0}^{\prime}.$
  2. (ii)
    A paired homomorphism is a paired map which is a homomorphism (taken in its appropriate context, e.g., semiring pair or nd-semiring pair).
  3. (iii)
    The paired kernel congruence of a paired map $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is $({\operatorname{\Phi}}_{f},{{\operatorname{\Phi}}_{f}}_{0})$, where ${\operatorname{\Phi}}_{f}:=\{(b_{1},b_{2})\in\mathcal{A}\times\mathcal{A}:f(b_{1})=f(b_{2})\}$ and ${{\operatorname{\Phi}}_{f}}_{0}:=\{(b_{1},b_{2})\in{\operatorname{\Phi}}_{f}:f(b_{1})\in\mathcal{A}^{\prime}_{0}\}.$
Lemma 2.21.
For any $\mathcal{T}$-module homomorphism $f:\mathcal{A}\to\mathcal{A}^{\prime}$ where $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is a pair, there is a pair $(\mathcal{A},\mathcal{A}_{0})$ where $\mathcal{A}_{0}=\{b\in\mathcal{A}:f(b)\in\mathcal{A}_{0}^{\prime}\},$ and then $f$ is a paired homomorphism.
Proof.   If $a\in\mathcal{T}$ and $b\in\mathcal{A}_{0}$ then $f(ab)=f(a)f(b)\in\mathcal{A}_{0}^{\prime},$ so $ab\in\mathcal{A}_{0}.$ Then $f(\mathcal{A}_{0})\subseteq\mathcal{A}^{\prime}_{0},$ by definition.
Lemma 2.22.
Suppose that $\operatorname{\Phi}$} is a congruence of the pair $(\mathcal{A},\mathcal{A}_{0})$, and let ${\operatorname{\Phi}}_{0}=(\mathcal{A}_{0}\times\mathcal{A}_{0})\cap{\operatorname{\Phi}}$.
  1. (i)
    Then $(\mathcal{A}/{\operatorname{\Phi}},\mathcal{A}_{0}/{{\operatorname{\Phi}}_{0}})$ is a pair, with the “projection” paired homomorphism $\pi:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}/{\operatorname{\Phi}},\mathcal{A}_{0}/{{\operatorname{\Phi}}_{0}})$ given by $b\mapsto[b].$
  2. (ii)
    If $\mathcal{A}_{0}\subseteq\mathcal{A}_{0}^{\prime}\subseteq\mathcal{A},$ then the identity map induces a paired map $(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A},\mathcal{A}_{0}^{\prime})$, which we call an extension of the null part.
Proof.   (i) Standard for congruences; one just checks that ${\operatorname{\Phi}}_{0}$ is a congruence of $\mathcal{A}_{0},$ being the restriction of $\operatorname{\Phi}$}.
(ii) Immediate.
Proposition 2.23.
Any paired map $\varphi:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ is the composite of the projection $(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}/{\operatorname{\Phi}}_{\varphi},\mathcal{A}_{0}/{{\operatorname{\Phi}}_{\varphi}}_{0})$ and an extension of the null part (up to isomorphism).
Proof.   By Lemma 2.22, there is an induced paired map $\bar{\varphi}:(\mathcal{A}/{\operatorname{\Phi}}_{\varphi},\mathcal{A}_{0}/{{\operatorname{\Phi}}_{\varphi}}_{0})\to(\mathcal{A}^{\prime},\varphi(\mathcal{A}_{0}))$ whose paired kernel congruence is trivial, since $\varphi([b_{1}])=\varphi([b_{2}])$ if and only if $\varphi(b_{1})=\varphi(b_{2}),$ if and only if $b_{1}=b_{2}$; afterwards the only change is an extension of the null part.
MAJOR NOTE 2.24.
Proposition 2.23 shows how ideals fit into the the structure theory of pairs, via extensions of the null part.
Definition 2.25.
Suppose $\mathcal{I}$ is a $\mathcal{T}$-submodule of $\mathcal{A}.$ An $\mathcal{I}$-congruence pair $({\operatorname{\Phi}},{\operatorname{\Phi}}_{\mathcal{I}})$ on $\mathcal{A}$ is a congruence $\operatorname{\Phi}$} on $\mathcal{A}\times\mathcal{A},$ together with its restriction $\operatorname{\Phi}$$_{\mathcal{I}}$ to $\mathcal{I}\times\mathcal{I}.$
Remark 2.26.
  1. (i)
    In particular, any congruence $\operatorname{\Phi}$} can also be viewed as the $\mathcal{A}_{0}$-congruence pair $({\operatorname{\Phi}},{\operatorname{\Phi}}_{0})$.
  2. (ii)
    Any paired map $\varphi:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ yields, taking $\mathcal{I}=\varphi^{-1}(\mathcal{A}^{\prime}_{0})$, the $\mathcal{I}$-congruence pair $({\operatorname{\Phi}}_{\varphi},{{\operatorname{\Phi}}_{\varphi}}_{\mathcal{I}})$, where ${\operatorname{\Phi}}_{\varphi}:=\{(b_{1},b_{2}):\varphi(b_{1})=\varphi(b_{2})\}.$
2.1.3. Examples of pairs
Let us see how to put all of our previous examples into the context of semiring pairs and pre-surpassing relations. Notation is from §1.2.
  1. (i)
    Some miscellaneous examples:
    • $\mathbb{V}_{n}$ example 1.3 becomes a pair satisfying Property N, when we put $\mathcal{T}=\{1\}$ and $\mathcal{A}_{0}=\{\mathbf{0},n\}.$
    • At the other extreme, [4, Example 2.21] provides the semiring pair $(\mathcal{A},\mathcal{A}_{0})$ with $\mathcal{A}=\mathcal{T}_{0}\cup\{e\}$, $\mathcal{T}$ an arbitrary monoid, and $\mathcal{A}_{0}=\{\mathbf{0},e\},$ satisfying $a_{1}+a_{2}=e$ for all $a_{1},a_{2}\in\mathcal{T}.$ The pair $(\mathcal{A},\mathcal{A}_{0})$ is $\mathcal{A}_{0}$-bipotent and shallow, but we have little control over it, and we need to cope with this example when formulating theorems.
    • An idempotent $\mathcal{A}_{0}$-bipotent pair: Let $\mathcal{A}=\mathcal{T}_{0}\cup\{\infty\}$ where $\mathbf{0}$ is the additive identity and $a+a=a$, but $a+a^{\prime}=\infty$ for each $a\neq a^{\prime}\in\mathcal{T}$; then $\mathcal{A}_{0}=\{\mathbf{0},\infty\}.$
    • Suppose $\mathcal{A}=\mathbb{N}[\Lambda]$ in commuting indeterminates $\{\lambda_{i}:1\leq i<\infty\}$ over the monoid $\mathcal{T}$ generated by the $\lambda_{i},$ We identify all elements of $\mathcal{A}$ of support at least of order 3 with $e$. $(\mathcal{A},\mathcal{A}_{0})$ is not weakly metatangible since $\lambda_{1}+\lambda_{2}\notin\mathcal{T}+\mathcal{A}_{0}.$
  2. (ii)
    Supertropical pairs. Suppose $\mathcal{G}$ is an ordered abelian monoid with absorbing minimal element $\mathbf{0}_{\mathcal{G}}$, and $\mathcal{T}$ is a monoid, together with a monoid homomorphism $\mu:\mathcal{T}\to\mathcal{G}$. The supertropical semiring $T(\mathcal{T};\mathcal{G})$ of §1.3.1 is naturally a $\mathcal{T}$-module under the given action. $(T(\mathcal{T};\mathcal{G}),\mathcal{G})$ is the supertropical pair. It is shallow (as opposed to many other examples), and the identity map is a negation map.
    There is a projection $\mu:\mathcal{A}\to\mathcal{G}$, so the supertropical pair lies over the max-plus algebra of $\mathcal{G},$ and has been studied in considerable detail. The supertropical pair $(\mathcal{A},\mathcal{G})$ is $\mathcal{G}$-bipotent. One of our original objectives has been to see how generally the results about the supertropical pair can be formulated.
    The supertropical pair is of the first kind, graded by the Boolean monoid $(\mathbb{B},+)=\{0,1\}$; the 0-component is $\mathcal{T}$ and the $1$-component is $\mathcal{G}.$
  3. (iii)
    Any $\mathcal{T}$-semiring $\mathcal{A}$ gives rise to a doubled pair where $\hat{A}$ is as in Definition 1.18, and the null part $\hat{\mathcal{A}}_{0}$ is as in Remark 1.19. $(\hat{\mathcal{A}},\hat{\mathcal{A}}_{0})$ has the negation map $(-)(b_{1},b_{2})=(b_{2},b_{1}).$
  4. (iv)
    Tropicalized pairs. If $(\mathcal{G},+)$ is an ordered abelian semigroup and $(\mathcal{L},\mathcal{L}_{0})$ is a semiring (resp. nd-semiring) pair, then $(\mathcal{L}\rtimes\mathcal{G},L_{\mathbf{0}}\rtimes\mathcal{G})$ is a semiring (resp. nd-semiring) pair with underlying monoid $\mathcal{T}\times\mathcal{G}$, under the action $(a,g)(\ell,g^{\prime})=(a\ell,g+g^{\prime})$ for $a\in\mathcal{T},$ $\ell\in\mathcal{L},$ $g,g^{\prime}\in\mathcal{G}$. [^5] We could perform the same construction for $(\mathcal{L},\mathcal{L}_{0})=(\{\mathbf{1},\infty\},\{\infty\})$ as in Note 2.2, with $\mathbf{1}+\mathbf{1}=\infty$. In this case $(\mathcal{L}\rtimes\mathcal{G},(\mathcal{L}\rtimes\mathcal{G})_{\mathbf{0}})$ is isomorphic to the supertropical pair.
  5. (v)
    Suppose $(\mathcal{A},\mathcal{A}_{0})$ is a semiring pair.
    1. (a)
      The function pair. $(\mathcal{A}^{S},\mathcal{A}_{0}^{S})$ is a semiring pair with underlying monoid defined as the elements of $\mathcal{T}^{S}$ having support of order 1.
    2. (b)
      The monoid semiring pair. If $\mathcal{M}$ is a monoid then $(\mathcal{A}[\mathcal{M}],\mathcal{A}_{0}[\mathcal{M}])$ is a semiring pair over the monoid $\mathcal{T}\cdot\mathcal{M}=\{ay:a\in\mathcal{T},\ y\in\mathcal{M}\}$.
    In each case of (v), we have altered $\mathcal{T};$ this will be discussed more formally in §8.1. But then, even if $(\mathcal{A},\mathcal{A}_{0})$ is metatangible, $(\mathcal{A}^{S},\mathcal{A}_{0}^{S})$ and $(\mathcal{A}[S],\mathcal{A}_{0}[S])$ need not be weakly metatangible.
2.1.4. Hyperpairs
Any hypersemigroup $\mathcal{H}$ is naturally a $\mathcal{T}$-module, where $\mathcal{T}=\mathcal{H}$, so $\mathcal{A}:=\mathcal{P}(\mathcal{H})$ is a $\mathcal{T}$-module, with the action viewed elementwise, and $\iota:\mathcal{H}\to\mathcal{A}$ is given by $\iota(a)=\{a\}$. For any $\mathcal{T}$-submodule $S_{0}$ of $\mathcal{A}:=\mathcal{P}(\mathcal{H})$, we get a pair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0})$ where $\mathcal{P}(\mathcal{H})_{0}=\{S\subseteq\mathcal{H}:\mathbf{0}\in S\}$ obviously is closed under addition. The pair $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0}),$ called the hyperpair of $\mathcal{H},$ is uniquely negated when $\mathcal{H}$ is a hypergroup. [^6] In line with our generalization of negation map to Property N, one can weaken the condition of the hypernegative 1-1 so as not to be unique, but only requiring the set e=11e=1-1 to be unique. $(\mathcal{P}(\mathcal{H}),\mathcal{P}(\mathcal{H})_{0}),$ is weakly metatangible if and only $\mathcal{H}$ is “stringent” in the sense of [8]. $\mathcal{P}(\mathcal{H})$ is $\subseteq$-distributive in the sense of Definition 2.29.
Hyperpairs are very well behaved, as we see in the next two results.
Lemma 2.27.
The surpassing relation $\subseteq$ on a hyperpair satisfies fissure, and thus is strongly $\mathcal{T}$-reversible by Lemma 2.7.
Proof.   Suppose $a_{0}\preceq\boxplus_{i=1}^{t}a_{i}$ for $a_{i}\in\mathcal{T}$. Then $a_{0}\preceq a_{1}\boxplus(a_{2}\boxplus\dots\boxplus a_{t}),$ so there is some $a\in(a_{2}\boxplus\dots\boxplus a_{t})$ for which $a_{0}\in a_{1}\boxplus a,$ proving fissure.
“Reversibility” as defined in [4] is a special case of Lemma 2.27.
Lemma 2.28.
Every hyperpair of a hyperfield is an $\mathcal{A}_{0}$-domain.
Proof.   If $\mathbf{0}\notin S_{1},S_{2},$ then $\mathbf{0}\notin S_{1}S_{2}.$
2.2. Distributivity
Because multiplication in hyperpairs, such as of the phase hyperfield of Example 1.15, may fail to be distributive over addition, we define variants of distributivity.
Definition 2.29.
  1. (1)
    The analogous definitions from 2.14 hold for nd-semiring pairs.
  2. (2)
    $(\mathcal{A},\mathcal{A}_{0})$ is $\preceq$-distributive with respect to a pre-surpassing relation $\preceq$, if $\sum_{i,j}a_{i}a_{j}^{\prime}\preceq\sum_{i}a_{i}\sum_{j}a_{j}^{\prime}$ for all $a_{i},a_{j}^{\prime}\in\mathcal{T}.$
Remark 2.30.
Hyperpairs are $\subseteq$-distributive.
3. Roots of polynomials
Remark 3.1.
  1. (i)
    Any paired homomorphism $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ extends naturally to a paired homomorphism $(\mathcal{A}^{S},\mathcal{A}^{S}_{0})\to({\mathcal{A}^{\prime}}^{S},{\mathcal{A}^{\prime}}^{S}_{0})$, defined elementwise, as well as to a paired homomorphism $(\mathcal{A}[\mathcal{M}],\mathcal{A}[\mathcal{M}]_{0})\to({\mathcal{A}^{\prime}}[\mathcal{M}],{\mathcal{A}^{\prime}}[\mathcal{M}]_{0})$.
  2. (ii)
    For any ${\mathbf{a}}=(a_{i})\in\mathcal{A}^{(i)}$ there is a paired homomorphism
    $(\mathcal{A}[\Lambda],\mathcal{A}[\Lambda]_{0})\to(\mathcal{A}^{(I)},\mathcal{A}_{0}^{(I)})$
    where $\Lambda=\{\lambda_{i}:i\in I\},$ given by $\lambda_{i}\to a_{i},$ $\forall i\in I.$
  3. (iii)
    (Special case for $|I|=1$) For any $a\in\mathcal{A}$ there is a paired homomorphism $(\mathcal{A}[\lambda],\mathcal{A}_{0}[\lambda])\to(\mathcal{A},\mathcal{A}_{0})$ given by $\lambda\to a,$ so $f\mapsto f(a)$ for all $f\in\mathcal{A}[\lambda].$
When $(\mathcal{A},\mathcal{A}_{0})$ is a semiring pair, these observations pertain to paired semiring homomorphisms.
MAJOR NOTE 3.2.
One basic tenet of classical algebra is that $a$ is a root of $f$ if and only if $f\mapsto 0$ under the homomorphism $\mathcal{A}[\lambda]\to\mathcal{A}$ sending $\lambda\mapsto a.$ But for pairs, this must be modified to make sense for our main examples. To start with, $f(a)=\mathbf{0}$ cannot hold for $f\neq\mathbf{0}$ in a ZSF semiring [^7] The way this is circumvented for semirings in general is to define a root of a pair $(f,g)$ of polynomials to be some $b\in\mathcal{A}$ such that $f(b)=g(b).$ But then one has to match roots with congruences, and many transparent results from algebra become contorted..
Accordingly we have a different notion of root.
Definition 3.3.
Suppose $\Lambda=\{\lambda_{i}:i\in I\}.$ An element ${\mathbf{b}}\in\mathcal{A}^{(I)}$ is a (null) root of $f(\Lambda)$ if $f({\mathbf{b}})\in\mathcal{A}_{0}^{(I)},$ i.e., ${\mathbf{b}}\in f^{-1}(\mathcal{A}_{0}^{(I)})$.
Remark 3.4.
If $b$ is a root of $f,$ then the paired homomorphism $(\mathcal{A},\mathcal{A}_{0})[\lambda]\to(\mathcal{A},\mathcal{A}_{0})[b]$ sending $\lambda\mapsto b$ sends $f$ to an element of $\mathcal{A}_{0}.$
Our motivation for this definition is the Cayley-Hamilton theorem 5.4 below.
In classical algebra, an element $b$ is a root of $f$ if and only if $\lambda-b$ divides $f.$ Thus, we want an appropriate notion of divisibility. The straightforward definition is not useful in general for polynomials and matrices, so we turn to alternatives.
3.1. Factorization with respect to roots of polynomials
Throughout this subsection, $(\mathcal{A},\mathcal{A}_{0})$ is a semiring pair, and $(\mathcal{A}[\lambda],\mathcal{A}[\lambda]_{0})$ is the polynomial pair in the indeterminate $\lambda.$ We modify the treatment in [18].
Definition 3.5.
There are three kinds of factorization, for $f_{1},f_{2}$ in $(\mathcal{A}^{S},\mathcal{A}^{S}_{0})$ or in $(\mathcal{A}[\Lambda],\mathcal{A}[\Lambda]_{0})$.
  • Usual division of elements in a monoid: $f_{1}|f_{2},$ if there is tangible $g$ such that $f_{2}=gf_{1}.$
  • $f_{1}$ $\preceq$-divides $f_{2}$, written $f_{1}\,|_{\preceq}\,f_{2},$ if $f_{2}\preceq gf_{1}$ for some tangible $g$.
  • $f_{1}$ $\mathcal{A}_{0}$-divides $f_{2},$ written $f_{1}\ |_{0}\ f_{2},$ if $f_{2}+gf_{1}\in\mathcal{A}_{0}$ for some tangible $g$.
(When $g$ is not required to be tangible, the definitions become vacuous.)
Remark 3.6.
If $f_{1}|_{\preceq}f_{2},$ then $f_{1}\ |_{0}\ f_{2}.$
Although division is antisymmetric up to multiplication by a unit in a cancellative monoid, the situation with is more intricate for polynomials, since the product of tangible polynomials need not be tangible. So far we have not been able to circumvent this difficulty to obtain a good theory of arithmetic. Nevertheless, there is a decent theory of roots of polynomials.
Example 3.7.
$\lambda+b^{\dagger}\ |_{\preceq}\ \lambda^{2}+(b^{2})^{\dagger}$, because
$(\lambda+b^{\dagger})(\lambda+b)=\lambda^{2}+(b^{2})^{\dagger}+(b+b^{\dagger})\lambda\succeq_{0}\lambda^{2}+bb^{\dagger}.$
Lemma 3.8.
Suppose $(\lambda-a)\ |_{\preceq}\ f,$ for a tangible polynomial $f$ i.e., $f\preceq(\lambda-a)g.$ Then $\deg g=\deg f-1.$
Proof.   The leading coefficient of $f$ is tangible, so must be matched by the leading coefficient of $g$, with the powers of $\lambda$ matching.
Lemma 3.9.
Given $b\in\mathcal{A},$ let $g_{b,n}=\sum_{j=0}^{n-1}b_{n-j}\lambda^{j}.$
  1. (i)
    $(\lambda+b^{\dagger})g_{b,n}=\lambda^{n}+(b^{n})^{\dagger}+\sum_{j=1}^{n-1}(b^{j}+(b^{j})^{\dagger})\lambda^{n-j}$.
  2. (ii)
    $\lambda^{n}+(b^{n})^{\dagger}\preceq(\lambda+b^{\dagger})g_{b,n}$.
  3. (iii)
    Suppose $f(\lambda)=\sum_{i=0}^{n}\alpha_{i}\lambda^{i}\in\mathcal{A}[\lambda].$ Let $g=\sum_{j=1}^{n}\alpha_{j}g_{b,j}\in\mathcal{A}[\lambda].$ Then
    $(\lambda+b^{\dagger})g=f(\lambda)+f(b)^{\dagger}+\sum_{i=0}^{n}\sum_{j=1}^{i-1}\alpha_{i}(b^{j}+(b^{j})^{\dagger})\lambda^{n-j}.$
    Hence $f(\lambda)+f(b)^{\dagger}\preceq(\lambda+b^{\dagger})g$.
Proof.   (i) Direct computation.
(ii) By (i), $(\lambda+b^{\dagger})g_{b,n}=\lambda^{n}+(b^{n})^{\dagger}+\sum(b^{j}+(b^{j})^{\dagger})\lambda^{n-j}\succeq\lambda^{n}+(b^{n})^{\dagger}$.
(iii) $f(\lambda)+f(b)^{\dagger}=\sum\alpha_{i}\lambda^{i}+\sum\alpha_{i}(b^{i})^{\dagger}=\sum\alpha_{i}(\lambda^{i}+(b^{i})^{\dagger}),$ so
$(f(\lambda)+f(b)^{\dagger})+\sum_{i=0}^{n}\sum_{j=1}^{i-1}\alpha_{i}(b^{j}+(b^{j})^{\dagger})\lambda^{n-j}=\sum\alpha_{i}(\lambda+b^{\dagger})g_{i}=(\lambda+b^{\dagger})g.$
(iv) By (iii) and Remark 3.6.
$f(\lambda)\preceq(\lambda+a^{\dagger})g$ since $f(a)\in\mathcal{A}_{0}.$ So to conclude that $(\lambda+a^{\dagger})\ |_{\preceq}\ f$ we $g$ need to be tangible.
Proposition 3.10.
Suppose $(\mathcal{A},\mathcal{A}_{0})$ is metatangible, and $f$ is a tangible polynomial of minimal degree for which $a$ is a tangible root of $f.$ Then $(\lambda+a^{\dagger})\ |_{\preceq}\ f(\lambda)$.
Proof.   $(\lambda+a^{\dagger})g=f(\lambda)+f(a)^{\dagger}+\sum_{i=0}^{n}\sum_{j=1}^{i-1}\alpha_{i}(a^{j}+(a^{j})^{\dagger})\lambda^{n-j},$ by Lemma 3.9(iii), implying $f(\lambda)\preceq(\lambda+a^{\dagger})g$ since $f(a)\in\mathcal{A}_{0}.$ The only problem is whether $g$ is tangible, but if not, then some tangible expression in $a$ of degree $\leq n$ is not tangible, so lies in $\mathcal{A}_{0}$.
Remark 3.11.
$(\lambda+a^{\dagger})\ |_{0}\ f(\lambda)$, under the hypothesis of Proposition 3.10, The converse is false in general, since one can have $a+a(-)a\in\mathcal{A}_{0}$, for $a\in\mathcal{T},$ in which case $\lambda(-)a\,|_{0}\,\lambda,$ and $a$ is not a root of $\lambda.$
For a counterexample to the converse of Proposition 3.10, let $\mathcal{F}$ be the polynomial semiring $\mathbb{N}[\lambda]$, and $\mathcal{A}$ be the doubled semiring of $\mathcal{F}[x_{1},x_{2},x_{3}]$, the polynomial semiring in three indeterminates, and take
$\mathcal{A}_{0}=\operatorname{diag}+(x_{1}x_{2}+x_{1}x_{3}+x_{2}x_{3})\mathcal{A}+(\sum x_{i})\mathcal{A}.$
$f(\lambda)=\lambda^{3}(-)x_{1}x_{2}x_{3}.$
$(\lambda(-)x_{1})(\lambda(-)x_{2})(\lambda(-)x_{3})=\lambda^{3}(-)\sum x_{i}\lambda^{2}+(x_{1}x_{3}+x_{2}x_{3}+x_{1}x_{2})\lambda(-)x_{1}x_{2}x_{3}=f(-)\sum x_{i}\lambda^{2}+(x_{1}x_{3}+x_{2}x_{3}+x_{1}x_{2})\lambda,$ implying $(\lambda-a_{i})\ |_{\preceq}\ f(\lambda)$ for each $i.$ But $f(x_{1})=x_{1}(x_{1}^{2}(-)x_{2}x_{3})\notin\mathcal{A}_{0},$ since it has too few monomials.
Definition 3.12.
Suppose $(\mathcal{A},\mathcal{A}_{0})$ is a pair with a negation map ()(-) and a pre-surpassing relation $\preceq.$ An element $a\in\mathcal{T}$ is a $\preceq$-root of $f$ if $f\preceq(\lambda(-)a)g$ for some tangible polynomial $g.$
Gunn [Proposition 4.7][17] has an argument for roots being $\preceq$-roots in a certain instance. We modify his argument to provide a rather general fundamental result.
Theorem 3.13.
Suppose $(\mathcal{A},\mathcal{A}_{0})$ has a $\mathcal{T}$-reversible surpassing relation $\preceq$ (which by Lemma 2.5 implies $(\mathcal{A},\mathcal{A}_{0})$ is uniquely negated), and $f\in\mathcal{A}[\lambda]$ with $a\in\mathcal{T}$.
  1. (i)
    If $a$ is a $\preceq$-root of $f$, then $a$ is a root of $f$.
  2. (ii)
    If $\preceq$ satisfies fissure, then each tangible root of any tangible polynomial is a $\preceq$-root.
Proof.   (i) Suppose $f\,|_{\preceq}\,(\lambda(-)a)g=\lambda g(-)ag.$ Write $f=\sum\alpha_{i}\lambda^{i}$ and $g=\sum\beta_{i}\lambda^{i}.$ Matching coefficients we get $\alpha_{i}\preceq(-)a\beta_{i}+\beta_{i-1}$, $i\geq 1,$ implying $(-)\beta_{i-1}\preceq(-)\alpha_{i}(-)a\beta_{i}$ since $\beta_{i-1}$ is presumed tangible, so $\beta_{i-1}\preceq\alpha_{i}+a\beta_{i}$, and $\alpha_{0}\preceq(-)a\beta_{0},$ implying $\mathbf{0}\preceq\alpha_{0}+a\beta_{0}\preceq\alpha_{0}+a\alpha_{1}+a^{2}\beta_{1}\preceq\sum_{i=0}^{n-1}\alpha_{i}a^{i}+a^{n}\beta_{n}=\sum_{i=0}^{n-1}\alpha_{i}a^{i}+\alpha_{n}a^{n}=f(a).$
(ii) Reversing the argument of (i), following the proof of [17, Proposition 4.7] (where I think one needs to take $f$ tangible).
Corollary 3.14.
Suppose that $f$ is a tangible polynomial, $a\in\mathcal{T}$, and $(\mathcal{A},\mathcal{A}_{0})$ has a $\mathcal{T}$-reversible surpassing relation, and either $\preceq$ satisfies fissure, or $(\mathcal{A},\mathcal{A}_{0})$ is metatangible and $a$ is not a root of any tangible polynomial of degree <n.<n. Then the roots of $f$ are precisely the $\preceq$-roots of $f$.
Proof.   (i) $(\Rightarrow)$ (ii) by Theorem 3.13(ii) Lemma 3.9 and (iii).
(ii) $(\Rightarrow)$ (i) by Theorem 3.13(i).
We can treat several $\preceq$-roots at once if $(\mathcal{A},\mathcal{A}_{0})$ is $\mathcal{A}$-prime.
Proposition 3.15.
If $(\mathcal{A},\mathcal{A}_{0})$ is $\mathcal{A}$-prime, then a monic polynomial $f$ cannot have more than $n$ distinct tangible $\preceq$-roots. If $a_{1},\dots,a_{n}$ are distinct tangible $\preceq$-roots then $f\preceq\prod_{i=1}^{n}(\lambda(-)a_{i}).$
Proof.   Writing $f\preceq(\lambda(-)a_{1})g$ for $g$ monic, we have $\mathbf{0}\leq f(a_{2})=(a_{2}(-)a_{1})g(a_{2}).$ By hypothesis, $g(a_{2})\in\mathcal{A}_{0},$ clearly of degree n1.n-1. Continuing inductively, and noting $\deg g=\deg f-1,$ we cannot have more than $n$ tangible $\preceq$-roots of $f$.
Corollary 3.16.
The Hamilton-Cayley polynomial of a matrix cannot have more than $n$ $\preceq$-roots.
Remark 3.17.
This leaves us with multiple roots. In classical algebra we say that $a$ is a double root of $f$ when $(\lambda(-)a)^{2}$ divides $f.$ Let us compare several options for generalization. Define the formal derivative $f^{\prime}$ of $f=\sum\alpha_{i}\lambda^{i}$ to be $\sum i\alpha_{i}\lambda^{i-1}.$
  1. (i)
    $a$ is a double $\preceq$-root of $f$ if $(\lambda(-)a)^{2}\,|_{\preceq}\,f,$ i.e., $(\lambda(-)a)^{2}h\succeq f$ for some tangible polynomial $h$.
  2. (ii)
    (Like Gunn’s version) $a$ is a double $\preceq$-root of $f$ if, for some tangible polynomial $g$, $f(-)(\lambda(-)a)g\in\mathcal{A}_{0}$ and $a$ is a $\preceq$-root of $g$.
  3. (iii)
    $a$ is a double $\preceq$-root of $f$ if $a$ is a $\preceq$-root of both $f$ and $f^{\prime}$.
These three versions are known to be equivalent in classical algebra, so which should we choose? Version (i) may seem the most natural at this stage, but we would like a double root also to be a single root. If $(\lambda(-)a)^{2}h\succeq f$ then clearly $(\lambda(-)a)g\succeq f$ for $g=(\lambda(-)a)h,$ but unfortunately $(\lambda(-)a)h$ need not be tangible. This leads to (ii), in which Descartes’ rule of signs has been proved for tropical extensions, in [6, 17, 5], but there is ambiguity with respect to $g$, which leads to complications. Therefore we turn to (iii), which is straightforward and unambiguous, but unfortunately $f^{\prime}$ need not be tangible! But $f^{\prime}$ indeed is tangible when no multiple of $a$ is in $\mathcal{A}_{0}$.
Note that $((\lambda(-)a)g)^{\prime}=g+(\lambda(-)a)g^{\prime}$; hence if $f\preceq(\lambda(-)a)g,$ then $f^{\prime}\preceq g+(\lambda(-)a)g^{\prime}$.
4. Geometry
Building on Definition 3.3(ii), given a set of polynomials $\mathcal{I}\subseteq\mathcal{A}[\Lambda],$ where $\Lambda=\{\lambda_{i}:i\in I\},$ define its $\mathcal{A}_{0}$-locus $Z(\mathcal{I})$ to be $\{{\mathbf{b}}\in\mathcal{A}^{(I)}:f({\mathbf{b}})\in\mathcal{A}_{0}^{(I)}\}$.
4.0.1. The Zariski correspondence
MAJOR NOTE 4.1.
There is a $\mathcal{T}$-semiring paired homomorphism
$\varphi_{\mathbf{b}}:(\mathcal{A}[\Lambda],\mathcal{A}[\Lambda]_{0})\to(\mathcal{A}^{(I)},\mathcal{A}^{(I))}_{0})$
defined by $\lambda_{i}\mapsto{\mathbf{b}}.$ Since, by Remark 2.26(ii), the paired homomorphism $(\varphi_{\mathbf{b}},\varphi_{{\mathbf{b}}_{0}}):(\mathcal{A}[\Lambda],\mathcal{A}[\Lambda]_{0})\to(\mathcal{A}^{(I)},\mathcal{A}^{(I))}_{0})$ is defined in terms of an $\mathcal{I}$-congruence pair ${\operatorname{\Phi}}_{\varphi_{\mathbf{b}}}$, where $\mathcal{I}=\varphi_{\mathbf{b}}^{-1}(\mathcal{A}^{\prime}_{0}),$ we can define the congruence ${\operatorname{cong}}(Z)=\cap_{{\mathbf{b}}\in Z}{\operatorname{\Phi}}_{\varphi_{\mathbf{b}}}.$
Conversely, any congruence $\operatorname{\Phi}$} on $(\mathcal{A}[\Lambda],\mathcal{A}[\Lambda]_{0})$ defines its $\mathcal{A}_{0}$-set $\{b\in\mathcal{A}:f(b)=g(b)\in\mathcal{A}_{0},\ \forall(f,g)\in{\operatorname{\Phi}}\}.$ Clearly $Z({\operatorname{cong}}(Z))=Z.$
Definition 4.2.
$(Z,{\operatorname{cong}})$ is the Zariski correspondence of pairs. A congruence $\operatorname{\Phi}$} of $(\mathcal{A}[\Lambda],\mathcal{A}[\Lambda]_{0})$ is radical if ${\operatorname{\Phi}}(Z({\operatorname{\Phi}}))=Z.$
Question 4.3.
Which congruences are radical (a Nullstellensatz)?.
4.1. Spectrum of prime congruences
Having identified $A_{0}$-loci with congruences, one is led to study the spectrum of prime congruences. Joo and Mincheva [30] came up with the following definition, which we state for arbitrary nd-semiring pairs.
Definition 4.4.
Suppose that $(\mathcal{A},\mathcal{A}_{0})$ is an nd-semiring pair.
  1. (i)
    The twist product ${\operatorname{\Phi}}_{1}\cdot_{\operatorname{tw}}{\operatorname{\Phi}}_{2}:=\{{\mathbf{b}}_{1}\cdot_{\operatorname{tw}}{\mathbf{b}}_{2}:{\mathbf{b}}_{i}\in{\operatorname{\Phi}}_{i}\}.$
  2. (ii)
    A congruence $\operatorname{\Phi}$} of $\mathcal{A}$ is semiprime if it satisfies the following condition: For a congruence ${\operatorname{\Phi}}_{1}\supseteq{\operatorname{\Phi}}$, if ${\operatorname{\Phi}}_{1}\cdot_{\operatorname{tw}}{\operatorname{\Phi}}_{1}\subseteq{\operatorname{\Phi}}$ then ${\operatorname{\Phi}}_{1}={\operatorname{\Phi}}$.
  3. (iii)
    A congruence $\operatorname{\Phi}$} of $\mathcal{A}$ is prime if it satisfies the following condition: For congruences ${\operatorname{\Phi}}_{1}$, ${\operatorname{\Phi}}_{2}\supseteq{\operatorname{\Phi}}$, if ${\operatorname{\Phi}}_{1}\cdot_{\operatorname{tw}}{\operatorname{\Phi}}_{2}\subseteq{\operatorname{\Phi}}$, then ${\operatorname{\Phi}}_{1}={\operatorname{\Phi}}$ or ${\operatorname{\Phi}}_{2}={\operatorname{\Phi}}$.
  4. (iv)
    A congruence $\operatorname{\Phi}$} of $(\mathcal{A},\mathcal{A}_{0})$ is irreducible if whenever ${\operatorname{\Phi}}_{1}\cap{\operatorname{\Phi}}_{2}={\operatorname{\Phi}}$ for congruences ${\operatorname{\Phi}}_{1}$, ${\operatorname{\Phi}}_{2}\supseteq{\operatorname{\Phi}}$, then ${\operatorname{\Phi}}_{1}={\operatorname{\Phi}}$ or ${\operatorname{\Phi}}_{2}={\operatorname{\Phi}}$.
Remark 4.5.
  • The intersection of semiprime congruences is semiprime.
  • A congruence $\operatorname{\Phi}$} is prime iff $\operatorname{\Phi}$} is semiprime and irreducible.
In classical algebra, Levitzki proved that any semiprime ideal $\operatorname{\Phi}$} is the intersection of prime ideals, which is a cornerstone in the theory of the prime spectrum.
Lemma 4.6.
(As in [52, Lemma 4.12]) Suppose $S\subseteq\widehat{A}$ is any subset disjoint from $\operatorname{\Phi}$}, satisfying the property that for any congruences $\Phi_{1},\Phi_{2}\supseteq\Phi,$ if $S\cap\Phi_{1},S\cap\Phi_{2}$ are nonempty then $S\cap(\Phi_{1}\cdot_{\operatorname{tw}}\Phi_{2})\neq\emptyset.$ Then there is a congruence ${\operatorname{\Phi}}^{\prime}\supseteq{\operatorname{\Phi}}$ of $\mathcal{A}$ maximal with respect to being disjoint from $S$, and $\operatorname{\Phi}$$^{\prime}$ is prime.
Proposition 4.7.
Any semiprime congruence $\operatorname{\Phi}$} of a semiring is the intersection of prime congruences.
Proof.   Following Levitsky, define a set $S\subset\widehat{A}$, by the following inductive method:
For any $s_{1}\in\widehat{A}\setminus{\operatorname{\Phi}},$ and given $s_{i}\in S$ take ${\mathbf{b}}\in\mathcal{A}$ and $s_{i+1}\in S$ such that $s_{i+1}\cdot_{\operatorname{tw}}{\mathbf{b}}\cdot_{\operatorname{tw}}s_{i+1}\notin{\operatorname{\Phi}}.$ By Lemma 4.6, we have some prime congruence disjoint $\operatorname{\Phi}$$^{\prime}$ from $S$. Hence the intersection of all such prime congruences misses all elements of $\widehat{A}\setminus{\operatorname{\Phi}},$ so is $\operatorname{\Phi}$}.
However, even in the idempotent commutative case, it is tricky to describe the intersection of all prime congruences, cf. [30, Example 3.3]. Joo and Mincheva prove that the prime congruence spectrum of $\mathbb{R}[\lambda_{1},\dots,\lambda_{n}]$ is catenary of length $n.$ This result is extended in [52] to semiring pairs satisfying the property $\mathbf{1}+ke=ke$ for some $k>0$. Although this property fails in classical algebra, it holds in the “miscellaneous examples” (i) of §2.1.3, supertropical pairs, all finite hyperpairs, and of course is preserved under extensions. There is much room for further investigation.
Question 4.8.
How far can one describe the prime spectrum? Is there an intrinsic description of the intersection of all prime congruences? Is there a “weak Nullstellensatz” describing semiprime congruences?
5. Matrix pairs
Definition 5.1.
The matrix pair $M_{n}(\mathcal{A},\mathcal{A}_{0})$ of a semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is $(M_{n}(\mathcal{A}),M_{n}(\mathcal{A}_{0}))$, which can be taken over the partial monoid $\cup_{1\leq i,j\leq n}\mathcal{T}e_{i,j}$.
5.1. Singularity of matrices
A track of an $n\times n$ matrix $A=(a_{i,j})$ is a product $a_{\pi}:=a_{\pi(1),1}\cdots a_{\pi(n),n}$ for $\pi\in S_{n}.$
We call a matrix tangible if all of its entries are in $\mathcal{T}.$ When $A$ is tangible, the tracks clearly exist and are in $\mathcal{T}_{0}$, so the following formulas make sense.
(5.1) ${\left|{A}\right|}_{+}=\sum_{\pi\in S_{n}\text{ even }}a_{\pi},\qquad{\left|{A}\right|}_{-}=\sum_{\pi\in S_{n}\text{ odd }}a_{\pi}.$
Also (5.1) makes sense over semiring pairs. This allows us to define singularity of matrices as follows:
Definition 5.2.
The matrix $A$ is singular [^8] In [4] one defines a balance relation $\nabla$ and singularity with respect to $\nabla$, but this definition is more encompassing. if ${\left|{A}\right|}_{+}={\left|{A}\right|}_{-}$.
5.2. $\dagger$}-determinants
Reutenauer and Straubing [47], and Zeilberger [59] showed that a number of determinantal identities of matrices admit bijective proofs, in the sense that the excess tracks pair off. This was the key idea in [4, §6], and is used again in the results given in this section and the next. The issue here is to avoid $(\mathbf{1}^{\dagger})^{2},$ since it may not be $\mathbf{1}.$
Suppose $\mathbf{1}+\mathbf{1}^{\dagger}=e$ in the pair $(\mathcal{A},\mathcal{A}_{0})$. We define the $\dagger$}-determinants $|A|_{\dagger}={\left|{A}\right|}_{+}+{\left|{A}\right|}_{-}^{\dagger}$ and $_{\dagger}|A|={\left|{A}\right|}_{+}^{\dagger}+{\left|{A}\right|}_{-}$. Familiar results can be reformulated for pairs:
Lemma 5.3.
For $A,B\in M_{n}(\mathcal{A})$,
  1. (i)
    $|AB|_{\dagger}\succeq_{0}|A|_{\dagger}|B|_{\dagger}$.
  2. (ii)
    $_{\dagger}|AB|\succeq_{0}{{}_{\dagger}}|A|{{}_{\dagger}}|B|$.
Proof.   The excess tracks on the left side pair off, contributing elements of $\mathcal{A}_{0}.$ [4, Theorem E] is generalized here to:
Theorem 5.4 (Cayley-Hamilton theorem).
For $A\in M_{n}(\mathcal{A}),$ viewed in $\widehat{\mathcal{A}}^{+}$ as in Remark 1.19 define the characteristic polynomial $f_{A}(\lambda)=|(\mathbf{0},\lambda I)+(A,\mathbf{0})|_{\dagger}.$ Then $f(A)\in\widehat{\mathcal{A}}_{0}.$
Proof.   By [4, Theorem E], applied to the doubled pair $(\hat{\mathcal{A}},\hat{\mathcal{A}}_{0}).$
5.3. $\dagger$}-adjoints
Definition 5.5.
Write $A_{i,j}$ for the $(j,i)$ minor of a matrix $A$. Define $a_{i,j}^{\prime}$ as $|A_{i,j}|_{\dagger}$ for $i+j$ even, and $_{\dagger}|A_{i,j}|$ for $i+j$ odd. The $\dagger$}-adjoint matrix $\operatorname{adj}_{\dagger}A$ is the matrix $(a_{i,j}^{\prime})$.
We reformulate [4, Theorem F].
Theorem 5.6 (Generalized Laplace identity and Cauchy-Binet formula).
$A=(a_{ij})$ satisfies Laplace’s well-known identity
$|A|=\sum_{j=1}^{n}a^{\prime}_{i,j}a_{i,j},$
for any $i$.
For the more general Cauchy-Binet formula, fix $I=\{i_{1},\dots,i_{m}\}\subset\{1,2,\dots,n\}.$ For any $J=\{j_{1},\dots,j_{m}\}\subset\{1,2,\dots,n\}$, write $A_{I,J}$ for the $(n-m)\times(n-m)$ minor obtained by deleting all rows from $I$ and all columns from $J$. Take $a_{I,J}^{\prime}$ for $|A_{I,J}|_{\dagger}$ when $(i_{1}+\dots+i_{m})(j_{1}+\dots+j_{m})$ is even, and for $_{\dagger}|A_{I,J}|$ when $(i_{1}+\dots+i_{m})(j_{1}+\dots+j_{m})$ odd. Then
(5.2) $|A|=\sum_{J:|J|=m}a^{\prime}_{I,J}|a_{I,J}|.$
Proof.   The proof goes over mutatus mutandus, since the formula was arrranged so that $\dagger$} does not repeat.
Definition 5.7.
A quasi-identity matrix is a matrix of the form $I+B_{0}$, where $B_{0}\in M_{n}(\mathcal{A}_{0})$ with $\mathbf{0}$ on the diagonal.
Corollary 5.8.
Theorem 5.6 immediately implies there are quasi-identity matrices $I_{A,{\dagger}},I_{{\dagger},A}$ such that $A\operatorname{adj}_{\dagger}A=|A|_{\dagger}\,I_{{\dagger},A}$ and $\operatorname{adj}_{\dagger}AA=\,_{\dagger}\!|A|\,I_{A,{\dagger}}.$
This notion is studied in depth in [27, 28], whose results generalize mutatus mutandus, using Property N.
5.4. Eigenvalues of matrices
We define eigenvalues by means of the surpassing relation $\preceq.$
Definition 5.9.
An eigenvector of a matrix $A$ is a vector $\mathbf{v}$} such that $A{\mathbf{v}}\succeq\alpha{\mathbf{b}}$; and $\alpha$ is called the corresponding eigenvalue.
Lemma 5.10.
If ${\mathbf{v}}_{j}$ are eigenvectors of $A$ with eigenvalues $\alpha_{j},$ and $B$ is the matrix whose $j$ column is ${\mathbf{v}}_{j},$ then $AB\succeq BD,$ where $D=\operatorname{diag}\{\alpha_{1},\dots,\alpha_{n}\}.$
Proof.   By definition of eigenvector.
Let us bring in Theorem 5.4.
Theorem 5.11.
The eigenvalues of $A$ are the roots of the characteristic polynomial $f_{A}(\lambda).$ If these roots are distinct, then $AB\succeq BD,$ where the columns of $B$ are $\mathcal{A}_{0}$-independent.
This coordinates with Corollary 3.16, when $\preceq$-roots are roots, and gives us a technique for diagonalization.
5.5. Matrix semigroups
The next natural step is to define the general linear monoid and special linear monoid. One might intuitively want $\mathrm{GL}(n,(\mathcal{A},\mathcal{A}_{0}))=\{A\in:{\left|{A}\right|}_{\dagger}\succeq a\}$ for some $a\in\mathcal{T}$, but this could be everything if $\mathcal{T}$ has values absorbed in addition by the entries of $\mathcal{A}$. We prefer:
Definition 5.12.
$\mathrm{GL}(n,(\mathcal{A},\mathcal{A}_{0}))=\{A\in M_{n}(\mathcal{A}):AB,BA$ are quasi-identity matrices for suitable $B\in M_{n}(\mathcal{A})\}$.
$\operatorname{SL}(n,(\mathcal{A},\mathcal{A}_{0}))=\{A\in\mathrm{GL}(n,(\mathcal{A},\mathcal{A}_{0})):{\left|{A}\right|}_{\dagger}\succeq\mathbf{1}\}$.
Lemma 5.13.
$\mathrm{GL}(n,(\mathcal{A},\mathcal{A}_{0}))$ and $\operatorname{SL}(n,(\mathcal{A},\mathcal{A}_{0}))$ are monoids.
We get monoid pairs by putting $\mathrm{GL}(n,(\mathcal{A},\mathcal{A}_{0}))_{0}$ (resp.  $\operatorname{SL}(n,(\mathcal{A},\mathcal{A}_{0}))_{0}$ be the singular matrices in $\mathrm{GL}(n,(\mathcal{A},\mathcal{A}_{0}))$ (resp.  $\operatorname{SL}(n,(\mathcal{A},\mathcal{A}_{0}))$).
5.5.1. Involutions
An anti-multiplicative map satisfies $f(b_{1}b_{2})=f(b_{2})f(b_{1}),$ $\forall b_{i}\in\mathcal{A}_{1}.$ An anti-morphism is a morphism which is an anti-multiplicative map. Thus we have three kinds: anti-homomorphisms, weak anti-morphisms, and $\preceq$-anti-morphisms.
Definition 5.14.
An involution on an nd-semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is an anti-morphism of degree 2, i.e., a map ()(*) satisfying ($\forall\ b,b_{i}\in\mathcal{A}$):
  1. (i)
    $(b^{*})^{*}=b$,
  2. (ii)
    $(b_{1}b_{2})^{*}=b_{2}^{*}b_{1}^{*},$
  3. (iii)
    $(b^{\dagger})^{*}=(b^{*})^{\dagger},$
  4. (iv)
    $\mathcal{T}^{*}=\mathcal{T}$,
  5. (v)
    (When $(\mathcal{A},\mathcal{A}_{0})$ has a given surpassing relation $\preceq$) if $b_{1}\preceq b_{2}$ in $\mathcal{A},$ then $b_{1}^{*}\preceq b_{2}^{*}.$
(v) is automatic for $\preceq_{\circ},$ since $(b_{1}+c^{\circ})^{*}=b_{1}^{*}+(c^{*})^{\circ}.$
Lemma 5.15.
$(c^{\circ})^{*}=(c^{*})^{\circ}.$
Proof.   $(c^{\circ})^{*}=(c(-)c)^{*}=c^{*}(-)c^{*}=(c^{*})^{\circ}.$
Example 5.16.
Examples of involutions on the matrix semialgebra $M_{n}({\mathcal{A}})$:
  1. (i)
    The transpose map on $M_{n}({\mathcal{A}})$ is an involution denoted by $A\mapsto A^{t}.$
  2. (ii)
    When $n=2m$, there is another involution, called the symplectic involution $(s)$, given by $\left(\begin{matrix}A_{11}&A_{12}^{\dagger}\\ A_{21}&A_{22}\end{matrix}\right)^{s}=\left(\begin{matrix}A_{22}^{t}&A_{12}^{t}\\ {A_{21}^{\dagger}}^{t}&A_{11}^{t}\end{matrix}\right)$, for $A_{ij}\in M_{m}(\mathcal{A}).$
MAJOR NOTE 5.17.
We can define $\preceq$-orthogonal matrices via the condition $(I,(\mathbf{0}))\preceq AA^{t},A^{t}A,$ and thereby define the $\preceq$-orthogonal monoids, special $\preceq$-orthogonal monoids, and $\preceq$-symplectic monoids. These would be special cases of $\preceq$-algebraic groups, defined in the natural way as varieties of pairs endowed with a quasi-inverse map $(gg^{-1}\succeq I)$, but these have not yet been investigated.
Definition 5.18.
A representation of a monoid is a homomorphism to $\mathrm{GL}(n,(\mathcal{A},\mathcal{A}_{0}))$.
Right multiplication by an element of a finite monoid can be viewed as a permutation matrix in the usual way. Since a permutation matrix has a single track, its determinant is tangible, in fact $\mathbf{1}$ or $\mathbf{1}^{\dagger}.$ However, to have a meaningful character theory, one needs to take traces.
When $A=(a_{i,j})$, write $\mathrm{tr}(A)$ for $\sum a_{i,i}.$
Remark 5.19.
$\mathrm{tr}(AB)=\mathrm{tr}(BA).$ Hence $\{A\in M_{n}(\mathcal{A},\mathcal{A}_{0}):\mathrm{tr}(A)\succeq\mathbf{0}\}$ produces a Lie pair [13].
Question 5.20.
How does one work out a character theory of monoid pairs?
6. Vector space pairs and linear algebra
Definition 6.1.
A vector space pair over a pair $(\mathcal{A},\mathcal{A}_{0})$ is a pair $(\mathcal{V},\mathcal{V}_{0}),$ where $\mathcal{V}:=\mathcal{A}^{(J)}$ and $\mathcal{V}_{0}=\mathcal{A}_{0}^{(J)}$. (Usually $J=\{1,\dots,n\}.$) The $\mathcal{T}$-bimodule operations are defined componentwise.
Different notions of dependence are given in [16, Chapter  19] and [1], but as one might expect, we focus on the ones arising from pairs and surpassing relations.
Lemma 6.2.
If $\mathbf{v_{n}}=\sum_{i=1}^{n-1}\mathbf{v_{i}},$ then the matrix $A$ whose rows are $\mathbf{v_{1}},\dots,\mathbf{v_{n}}$ is singular.
Proof.   The ++ and - parts in (5.1) match.
Definition 6.3.
A set of vectors $\{\mathbf{v}_{i}\in\mathcal{V}:i\in I\}$ is $\mathcal{V}_{0}$-dependent, if $\sum_{i\in I^{\prime}}a_{i}\mathbf{v}_{i}\in\mathcal{V}_{0}$ for some nonempty finite subset $I^{\prime}\subseteq I$ and $a_{i}\in\mathcal{T}$.
$\mathcal{V}_{0}$-dependence is studied in terms of matrices, as in classical linear algebra, via the following conditions.
  • Condition A1: The submatrix rank is less than or equal to the row rank and the column rank.
  • Condition A2: The submatrix rank is greater than or equal to the row rank and the column rank.
  • Condition A3: If the row rank and the column rank of an $n\times n$ matrix $A$ is $n$, then $A$ is nonsingular.
  • Condition A4, cf. [7, Question 5.13]: The rows of an $m\times n$ matrix are $\mathcal{V}_{0}$-dependent if $m>n.$
  • Condition A5: For $m>n,$ either the rows of an $m\times n$ matrix $A$ are $\mathcal{V}_{0}$-dependent or $A$ has a singular $n\times n$ submatrix.
Remark 6.4.
  1. (i)
    Each of these conditions hold for supertropical algebras [26].
  2. (ii)
    Each of these conditions has a counterexample in [4]. There are straightforard counterexamples to Conditions A2 and A3.
  3. (iii)
    Condition A1 holds in varied situations, including
    • matrices over a metatangible, uniquely negated pair [4, §7.3.1].
    • quotient hyperfields, by the argument in [7] of passing down from the field.
    • Various other hypotheses yielding Condition A1 are given in [4, §7.3 and §7.5].
  4. (iv)
    Condition A2 implies A3.
  5. (v)
    Condition A3 is seen to hold in certain cases, in [4, Theorems O,P,Q].
  6. (vi)
    Condition A3 implies A4, since adding on mnm-n columns of zeroes gives a singular matrix, whose rows must thus be $\mathcal{V}_{0}$-dependent.
  7. (vii)
    Condition A4 holds in many situations, cf. [4, §7.7], including all hyperfield examples given in [4, §7.7.1].
When $\mathcal{A}$ has a pre-surpassing relation, the pre-surpassing relation $\preceq$ on vectors in $V$ is defined componentwise, i.e., $(v_{j})\preceq(v_{j}^{\prime})$ if $v_{j}\preceq v_{j}^{\prime}$ for each $i$. Although $\mathcal{V}_{0}$-dependence is our predominant tool, the following notion also is useful.
Definition 6.5.
A vector $\mathbf{v}$ is $\preceq$-spanned by vectors $\mathbf{v}_{1},\dots,\mathbf{v}_{n}$ if $\mathbf{v}\preceq\sum_{i=1}^{n}a_{i}\mathbf{v}_{i}$ for suitable $a_{i}\in\mathcal{T}_{0}$.
Remark 6.6.
For a $\mathcal{T}$-reversible pre-surpassing relation $\preceq$, if $\mathbf{v}_{1},\dots,\mathbf{v}_{n}$ are tangible and $\mathbf{v}_{1}$ is $\preceq$-spanned by $\mathbf{v}_{2},\dots,\mathbf{v}_{n}$, then $\mathbf{v}_{2}$ is $\preceq$-spanned by $\mathbf{v}_{1},\mathbf{v}_{3},\dots,\mathbf{v}_{n}$.
When $(\mathcal{A},\mathcal{A}_{0})$ is shallow, Condition A5 holds by [4, Proposition 7.4]. In fact, by [4, Theorem U], either $A$ has a singular $n\times n$ submatrix or each row is $\preceq_{0}$-spanned by the others.
One of the main tools for solving linear equations is Cramer’s rule, which is related to Condition 1 and studied in depth in [4].
7. Varieties and polynomial identities
It is natural to study varieties of pairs, either from the point of view of congruences or noncommutative $\mathcal{A}_{0}$-loci, leading to the notion of polynomial identity:
Definition 7.1.
Suppose $f(X_{1},\dots,X_{n})$, $g(X_{1},\dots,X_{n})$ are tangible polynomials in noncommuting indeterminates over a semiring pair $(\mathcal{A},\mathcal{A}_{0})$.
  • $f(X_{1},\dots,X_{n})=g(X_{1},\dots,X_{n})$ is a polynomial identity (PI) of a semiring $\mathcal{A}$ if $f(b_{1},\dots,b_{n})=g(b_{1},\dots,b_{n})$ for all $b_{i}\in\mathcal{A};$
  • $f(X_{1},\dots,X_{n})$ is an $\mathcal{A}_{0}$-PI of $(\mathcal{A},\mathcal{A}_{0})$ if $f(b_{1},\dots,b_{n})\in\mathcal{A}_{0}$ for all $b_{i}\in\mathcal{A}.$
Remark 7.2.
For any semiring pair $(\mathcal{A},\mathcal{A}_{0})$, if $f=g$ is a PI of $\mathcal{A}$, then $f+g^{\dagger}$ is an $\mathcal{A}_{0}$-PI of $(\mathcal{A},\mathcal{A}_{0})$.
By viewing any commutative semiring as a homomorphic image of a free commutative semiring $\mathbb{N}[\Lambda]$, which in term is contained in the free commutative ring $\mathbb{Z}[\Lambda]$, Akian, Gaubert and Guterman [1] observed that any polynomial expression holding in all commutative rings also holds in all commutative semirings. By passing to matrix semirings, and matching components, sorting positive and negative parts, it follows that all PI’s of $M_{n}(\mathbb{Z})$ yield PI’s of $M_{n}(\mathcal{A})$ for any commutative semiring $\mathcal{A},$ and thus $\mathcal{A}_{0}$-PIs of $M_{n}(\mathcal{A},\mathcal{A}_{0}).$ [1] also contains a “strong transfer principle,” building on Reutenauer and Straubing [47]. This raises the “Specht-type” question,
Question 7.3.
Do all $\mathcal{A}_{0}$-PIs of $M_{n}(\mathcal{A},\mathcal{A}_{0})$ follow formally from finitely many $\mathcal{A}_{0}$-PIs?
8. Categories: Morphisms of pairs
So far the only morphisms we have discussed are homomorphisms in the sense of universal algebra. A subtle categorical issue is to determine the most useful definition of morphism of pairs for applications. The theory flows most easily when we use the universal algebra approach with paired homomorphisms, as seen in [53]. Pairs are closed under products and sums in universal algebra.
However, there are two other definitions of morphisms which cast better light on many non-classical examples.
Definition 8.1.
All maps $f:(\mathcal{A},\mathcal{A}_{0})\to(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$ are presumed to satisfy $f(\mathcal{A}_{0})\subseteq\mathcal{A}_{0}^{\prime}$ and $f(\mathbf{0})=\mathbf{0}^{\prime}.$ We distinguish by how $f$ acts additively.
  1. (1)
    $f$ is a weak morphism of pairs if, for all $n\in\mathbb{N},$ $\sum_{i=1}^{n}b_{i}\in\mathcal{A}_{0}$ implies $\sum_{i=1}^{n}f(b_{i})\in\mathcal{A}^{\prime}_{0}$, for $b_{i}\in\mathcal{A}$.
  2. (2)
    $f$ is a $\preceq$-morphism with respect to pre-surpassing relations $\preceq$ of $(\mathcal{A},\mathcal{A}_{0})$ and $(\mathcal{A}^{\prime},\mathcal{A}^{\prime}_{0})$, if $f(b)\preceq f(b^{\prime})$ for $b\preceq b^{\prime}$ and $f(\sum_{i=1}^{n}b_{i})\preceq\sum_{i=1}^{n}f(b_{i})$ for all $n\in\mathbb{N},$ $b_{i}\in\mathcal{A}$.
  3. (3)
    A weak morphism (resp.  $\preceq$-morphism) of semiring pairs (or, more generally, nd-semiring pairs) is a weak morphism (resp. $\preceq$-morphism) which also is a multiplicative map.
  4. (4)
    $\succeq$-morphisms are defined analogously.
Lemma 8.2.
(As in [3, Lemma 2.10]) Every $\preceq$-morphism or $\succeq$-morphism of a pair with respect to a pre-surpassing relation is a weak morphism.
Remark 8.3.
The $\dagger$}-determinant on matrices is a $\succeq_{0}$-morphism, by Lemma 5.3.
Example 8.4.
For any pair $(\mathcal{A},\mathcal{A}_{0})$ and any $a\in\mathcal{T},$ the left multiplication map $y\mapsto ay$ is a paired homomorphism. When $(\mathcal{A},\mathcal{A}_{0})$ is a $\preceq$-distributive nd-semiring pair, $L_{b}$ is a $\preceq$-homomorphism. When $(\mathcal{A},\mathcal{A}_{0})$ is a semiring pair, the left multiplication map $L_{b}:y\mapsto by$ is a paired homomorphism for all $b\in\mathcal{A}$.
Lemma 8.5.
Take $\mathcal{T}=\mathcal{G}$ in §1.3.1, with $\operatorname{val}:\mathbb{K}\to\mathcal{T}$ the Puiseux valuation. Then the composition $\operatorname{val}:\mathbb{K}\to\mathcal{T}$ yields a $\preceq$-morphism $(\mathbb{K},\{0\})\to\mathbb{(}T(\mathcal{T};\mathcal{G}),\mathcal{G})$.
Proof.   $\operatorname{val}(f+g)=\max\{-\operatorname{val}f,-\operatorname{val}g\}$, which is the sum in $\mathbb{T}(\mathcal{T};\mathcal{G}),$ unless $\operatorname{val}f=\operatorname{val}g$, in which case $\operatorname{val}(f+g)\leq-\operatorname{val}f\preceq_{0}\mu(-\operatorname{val}f),$ which is the sum of $\operatorname{val}f$ and $\operatorname{val}f$ in $\mathbb{T}(\mathcal{T};\mathcal{G})$.
Remark 8.6.
Let $(\mathcal{V},\mathcal{V}_{0})$ be a vector space pair.
  1. (i)
    The left multiplication map by $a\in\mathcal{T}$ preserves dependence of vectors if and only if it is a weak morphism of $(\mathcal{V},\mathcal{V}_{0})$.
  2. (ii)
    The left multiplication map by $a\in\mathcal{T}$ preserves $\preceq$-dependence of vectors if and only if it is a $\preceq$-morphism of $(\mathcal{V},\mathcal{V}_{0})$.
The terminology for morphisms in hyperfields vary (and find their origin in [12]). Here we say a map $f:\mathcal{H}\to\mathcal{H}^{\prime}$ of hypergroups is a weak morphism if $0\in\boxplus a_{i}$ implies $0\in\boxplus f(a_{i})$; $f$ is a strong morphism if $f(\boxplus a_{i})\subseteq\boxplus f(a_{i}).$
Lemma 8.7.
There is a functor from hypergroups to hyperpairs, sending a map $f:\mathcal{H}\to\mathcal{H}^{\prime}$ of hypergroups to the map $\hat{f}:(\mathcal{P}^{\sharp}(\mathcal{H}),\mathcal{P}^{\sharp}(\mathcal{H}_{0}))\to(\mathcal{P}^{\sharp}(\mathcal{H}^{\prime}),\mathcal{P}^{\sharp}(\mathcal{H}^{\prime}_{0}))$ given by $\hat{f}(S)=\{f(a):a\in S\}.$
  1. (i)
    If $f$ is a weak morphism of hypergroups then $\hat{f}$ is a weak morphism of hyperpairs.
  2. (ii)
    If $f$ is a strong morphism of hypergroups then $\hat{f}$ is a $\subseteq$-morphism .
Proof.   Clearly if $S_{1}\subseteq S_{2}$ then $f(S_{1})\subseteq f(S_{2}),$ and
(8.1) $\hat{f}(S_{1}\boxplus S_{2})$ $=\{f(a):a\in S_{1}\boxplus S_{2}\}=\{f(a_{1}\boxplus a_{2}):a_{i}\in S_{i}\}$
$\subseteq\{f(a_{1})\boxplus f(a_{2}):a_{i}\in S_{i}\}=\hat{f}(S_{1})\boxplus\hat{f}(S_{2}).$
If $\mathbf{0}\in S$ then $\mathbf{0}=f(\mathbf{0})\in f(S).$
8.1. Extensions of pairs
So far, the underlying monoid $\mathcal{T}$ has been fixed in our discussion of pairs. But the polynomial and matrix constructions motivate changing the underlying monoid.
Definition 8.8.
$(\mathcal{A},\mathcal{A}_{0})$ is a subpair of a $\mathcal{T}^{\prime}$-pair $(\mathcal{A}^{\prime},\mathcal{A}_{0}^{\prime})$ is a $\mathcal{T}$-pair $(\mathcal{A},\mathcal{A}_{0})$ such that $\mathcal{A}\subseteq\mathcal{A}^{\prime},$ $\mathcal{T}=\mathcal{T}^{\prime}\cap\mathcal{A},$ and $\mathcal{A}_{0}=\mathcal{A}^{\prime}_{0}\cap\mathcal{A}.$
An extension of a pair $(\mathcal{A},\mathcal{A}_{0})$ is a $\mathcal{T}^{\prime}$-pair $(\mathcal{A}^{\prime},\mathcal{A}_{0}^{\prime})$ such that $(\mathcal{A},\mathcal{A}_{0})$ is a subpair of $(\mathcal{A}^{\prime},\mathcal{A}_{0}^{\prime}).$
The “tropical extension” of Definition 1.6 is an example. Here are some more.
8.1.1. Tensor extensions
When $\mathcal{A}$ is a $\mathcal{T}$-module, and $\mathcal{T}^{\prime}$ is an abelian monoid containing $\mathcal{T},$ we can define the tensor extension $\mathcal{T}^{\prime}$-module $\mathcal{A}\otimes_{\mathcal{T}}\mathcal{T}^{\prime}$ to be the $\mathcal{T}^{\prime}$-module $\mathcal{S}(\mathcal{T}^{\prime}\times\mathcal{A})/{\operatorname{\Phi}},$ where $\operatorname{\Phi}$} is the congruence generated by all
(8.2) $\bigg((ax,v),(x,av)\bigg),$
(8.3) $\bigg(\big(x,v+w\big),\big(x,v)+(x,w\big)\bigg),\qquad\forall x\in\mathcal{T}^{\prime},\,v,w\in\mathcal{\mathcal{A}}_{i},\,a\in\mathcal{T},$
cf. [53, §4.3]. When $(\mathcal{A},\mathcal{A}_{0})$ is a pair, $(\mathcal{A}\otimes_{\mathcal{T}}\mathcal{T}^{\prime},\mathcal{A}_{0}\otimes_{\mathcal{T}}\mathcal{T}^{\prime})$ is a $\mathcal{T}^{\prime}$-pair, called the $\mathcal{T}^{\prime}$-tensor pair extension. Homomorphisms, weak morphisms and $\preceq$-morphisms all lift up tensor pair extensions, by [53, Theorem 5.15].
Example 8.9.
Tensor pair extensions relate to our previous constructions of extensions.
  1. (i)
    $(\mathcal{A}[\lambda],\mathcal{A}_{0}[\lambda])\cong(\mathcal{A},\mathcal{A}_{0})\otimes\mathcal{T}^{\prime}$ where $\mathcal{T}^{\prime}=\{a_{i}\lambda^{i}:i\in\mathbb{N}\}.$
  2. (ii)
    $(\mathcal{A}^{S},\mathcal{A}_{0}^{S})\cong(\mathcal{A},\mathcal{A}_{0})\otimes\mathcal{T}^{\prime}$ where $\mathcal{T}^{\prime}$ is the submonoid of elements of $\mathcal{T}^{S}$ having support of order 1.
  3. (iii)
    If $\mathcal{T}$ is contained in a group $\mathcal{T}^{\prime}$ then $(\mathcal{A},\mathcal{A}_{0})\otimes\mathcal{T}^{\prime}$ is a pair over the group $\mathcal{T}^{\prime}$. This is an easy instance of localization.
8.1.2. Algebraic and integral extensions
Definition 8.10.
Suppose $(\mathcal{A}^{\prime},\mathcal{A}_{0}^{\prime})$ is an extension of $(\mathcal{A},\mathcal{A}_{0})$.
  • An element $b\in\mathcal{A}^{\prime}$ is $\preceq$-algebraic over $\mathcal{A}$ if $y\preceq f(b)$ for some $y\in\mathcal{A}$ and some tangible polynomial $f$; $b$ is $\preceq$-integral if $f$ is monic.
  • $(\mathcal{A}^{\prime},\mathcal{A}_{0}^{\prime})$ is $\preceq$-algebraic (resp.  $\preceq$-integral) over $(\mathcal{A},\mathcal{A}_{0})$ if each element of $\mathcal{A}^{\prime}$ is algebraic (resp.  integral) over $\mathcal{A}.$
Polynomial extensions are not algebraic.
Question 8.11.
What is a good general theory of integral extensions?
9. Module pairs
Motivated by the study of rings, we would like to use modules (often called “semimodules” in the literature) analogously in the study of semiring pairs. The theory of projective [34] and injective modules over semirings is carried out in the usual categorical setting using homomorphisms, also cf. [16, Chapter  17]. This has led to the search for a homology theory [54, 35].
Definition 9.1.
Assume that $\mathcal{A}$ is a $\mathcal{T}$-semiring, and that $\mathcal{M}$ is a $\mathcal{T}$-module satisfying $(ay)b=a(yb)=y(ab)$ for all $a\in\mathcal{T},$ $b\in\mathcal{A},$ $y\in\mathcal{M}.$ Also assume that there is an action $\mathcal{A}\times\mathcal{M}$, which restricts to the given $\mathcal{T}$-action.
  1. (i)
    $\mathcal{M}$ is a module over $\mathcal{A}$ if $\sum_{i,j}b_{i}y_{j}=(\sum_{i}b_{i})(\sum_{j}y_{j})$, for all $b_{i}\in\mathcal{A},$ $y_{j}\in\mathcal{M}.$
  2. (ii)
    $\mathcal{M}$ is a $\preceq$-module over $\mathcal{A}$ with respect to a surpassing relation $\preceq$ if $\sum_{i,j}b_{i}y_{j}\preceq(\sum_{i}b_{i})(\sum_{j}y_{j})$, for all $b_{i}\in\mathcal{A},$ $y_{j}\in\mathcal{M}.$
  3. (iii)
    A module pair over a semiring pair $(\mathcal{A},\mathcal{A}_{0})$ is a pair $(\mathcal{M},\mathcal{M}_{0})$ such that $\mathcal{M}\supseteq\mathcal{M}_{0}$ are modules over $\mathcal{A}$ and $\mathcal{A}_{0}\mathcal{M}\subseteq\mathcal{M}_{0}.$ $\preceq$-module pairs are defined analogously.
Remark 9.2.
Any hyperpair is a $\subseteq$-module pair over itself.
9.0.1. Tensor products of module pairs
Next, we take a special case of [53], which considers left and right $\mathcal{T}$-modules.
Definition 9.3.
  1. (i)
    Suppose $\mathcal{A}_{1},\mathcal{A}_{2}$ are $\mathcal{T}$-modules. Define the tensor product $\mathcal{T}$-module $\mathcal{A}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{2}$ to be the semigroup $\mathcal{S}(\mathcal{A}_{1}\times\mathcal{A}_{2})/{\operatorname{\Phi}},$ where $\operatorname{\Phi}$} is the congruence generated by all
    (9.1) $\bigg((ax_{1},x_{2}),(x_{1},ax_{2})\bigg),$
    (9.2) $\bigg(\big(v_{1}+w_{1},x_{2}\big),\big(v_{1},x_{2})+(w_{1},x_{2}\big)\bigg),$
    (9.3) $\bigg(\big(x_{1},v_{2}+w_{2}\big),(x_{1},v_{2})+(x_{1},w_{2})\bigg),$
    $\forall x_{i},v_{i},w_{i}\in\mathcal{\mathcal{A}}_{i},$ $a\in\mathcal{T}$.
  2. (ii)
    When $(\mathcal{A}_{i},{\mathcal{A}_{i}}_{0})$ are pairs, define
    $(\mathcal{A}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{2})_{0}:=({\mathcal{A}_{1}}_{0}\otimes_{\mathcal{T}}\mathcal{A}_{2})+(\mathcal{A}_{1}\otimes_{\mathcal{T}}{\mathcal{A}_{2}}_{0}).$
$((\mathcal{A}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{2}),(\mathcal{A}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{2})_{0})$ is a pair, called the tensor pair.
By [53, Corollaries [4.13,4.15], the tensor product satisfies the usual associativity and distributivity over direct sums. There is a glitch in defining multiplication on the tensor product $\mathcal{A}_{1}\otimes_{\mathcal{T}}\mathcal{A}_{2}$ of $\mathcal{T}$-semirings, since $\sum a_{i,1}\otimes\sum a_{i,2}$ and $\sum(a_{i,1}\otimes a_{i,2})$ may differ. We can avoid this difficulty if say $\mathcal{A}_{1}$ is free over $\mathcal{T}.$
Tensor products of module pairs are seen in [53, Corollaries 4.15, 4.16] to provide a concrete tensor category with respect to homomorphisms, in the sense of [29, 46], and satisfy the adjoint correspondence [53, Remark 5.17]. Tensor products of weak morphisms and $\preceq$-morphisms are considerably subtler. treated in Tensor products of module pairs are seen in [53, Corollaries 4.15, 4.16] to provide a concrete tensor category with respect to homomorphisms, in the sense of [29, 46], and satisfy the adjoint correspondence [53, Remark 5.17]. Tensor products of weak morphisms and $\preceq$-morphisms are considerably subtler, treated in [53, §5].
9.1. $\preceq$-Projective and $\preceq$-injective modules
Definition 9.4 ([31]).
  1. (i)
    A $\preceq$-morphism $f:(\mathcal{M},\mathcal{M}_{0})\to(\mathcal{N},\mathcal{N}_{0})$ is $\preceq$-onto if for every $b\in N$, there exists $a\in M$ for which $b\preceq f(a)$.
  2. (ii)
    A $\preceq$-module pair $(\mathcal{P},\mathcal{P}_{0})$ is $\preceq$-projective if for any $\preceq$-onto $\preceq$-morphism $h:(\mathcal{M},\mathcal{M}_{0})\to(\mathcal{N},\mathcal{N}_{0}),$ every $\preceq$-morphism $f:(\mathcal{P},\mathcal{P}_{0})\to(\mathcal{M},\mathcal{M}_{0})$ lifts to a $\preceq$-morphism $\tilde{f}:(\mathcal{P},\mathcal{P}_{0})\to(\mathcal{M},\mathcal{M}_{0})$, in the sense that $f=h\tilde{f}$.
The $\succeq$-version is analogous. The $\preceq$-projective theory is developed in [31], to include a “dual basis lemma” and a “Shnauel’s lemma.” In [32] this is carried forward towards homological theory with mixed success: A partial “connecting lemma” is obtained in [32, Theorem 5.15]. Much work remains to obtain a useable $\preceq$-homology theory.
Localization via injective modules is treated in [16, Chapter 18]. One can dualize Definition 9.4, to obtain $\preceq$-injective modules and the $\preceq$-injective hull.
Question 9.5.
How does one develop a $\preceq$-quotient theory?